### v1

# SubTrack your Grad: Gradient Subspace Tracking for Memory and Time Efficient Full-Parameter LLM Training

Abstract
Training Large Language Models (LLMs) demand significant time and computational resources due to their large model sizes and optimizer states. To overcome these challenges, recent methods, such as BAdam, employ partial weight updates to enhance time and memory efficiency, though sometimes at the cost of performance. Others, like GaLore, focus on maintaining performance while optimizing memory usage through full parameter training, but may incur higher time complexity. By leveraging the low-rank structure of the gradient and the Grassmannian geometry, we propose SubTrack-Grad, a subspace tracking-based optimization method that efficiently tracks the evolving gradient subspace by incorporating estimation errors and previously identified subspaces. SubTrack-Grad delivers better or on-par results compared to GaLore, while significantly outperforming BAdam, which, despite being time-efficient, compromises performance. SubTrack-Grad reduces wall-time by up to 20.57% on GLUE tasks (15% average reduction) and up to 65% on SuperGLUE tasks (22% average reduction) compared to GaLore. Notably, for a 3B parameter model, GaLore incurred a substantial 157% increase in wall-time compared to full-rank training, whereas SubTrack-Grad exhibited a 31% increase—representing a 49% reduction in wall-time, while enjoying the same memory reductions as GaLore.
Machine Learning, ICML
Figure 1: Visualization of SubTrack-Grad: Between subspace updates, gradients are projected onto a fixed subspace. The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error. The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
1 Introduction
Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity. However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact due to a larger carbon footprint. (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024). Consequently, there is a pressing need to develop time- and memory-efficient methods to make LLMs more accessible and reduce their environmental impact, while preseving performance.
Several techniques have been proposed to address memory challenges, including gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020). In this context, several approaches aim to reduce memory usage by optimizing a subset of model parameters or operating within a lower-dimensional space (Dettmers et al., 2024; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024; Hu et al., 2021). Notably, the well-known method LoRA (Hu et al., 2021), addresses memory efficiency by decomposing weight matrices into two low-rank trainable matrices, enabling the optimization of network parameters within a smaller subspace and significantly reducing memory usage. However, LoRA’s effectiveness relies on the assumption that parameters reside in a low-dimensional space, which does not always hold true. On the other hand, recent methods including BAdam (Luo et al., 2024) and the one introduced by Ramesh et al. (2024) leverage the block coordinate descent (BCD) framework and partition the model into different blocks to optimize one block at a time. These methods can effectively reduce memory footprint while avoiding extra computation; however, while these achieve better memory and time efficiency, they suffer from performance degradation due to partial parameter tuning.
Memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024). To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024). Prominantly, GaLore (Zhao et al., 2024) reduces the memory consumption of optimizers by projecting gradient matrices into a low-rank subspace and tracking changes through periodic singular value decomposition (SVD) to obtain a rank-$r$ approximation. Unlike weight parameters, there are many works showing that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023); using this fact, GaLore decreases optimizer memory footprint while enabling full and simultaneous parameter updates. However, their approach faces several challenges. First, SVD is computationally expensive, and if the gradient does not evolve within a nearly constant subspace, GaLore must increase the frequency of SVD operations, significantly increasing the amount of computation. This is problematic because not all layers’ gradients converge to a stable subspace early in training (Jaiswal et al., 2024). Moreover, applying SVD to a single gradient matrix is susceptible to data noise (Vaswani et al., 2018), and GaLore does not leverage 1) information from the orthogonal space as feedback to adjust the subspace (Modoranu et al., 2024) or 2) previously computed subspaces to incorporate past knowledge, which could help mitigate noise effects and improve convergence.
To this end, we propose Gradient Subspace Tracking (SubTrack-Grad), a Grassmannian-based, time- and memory-efficient method that adjusts the subspace using rank-$1$ updates for full parameter fine-tuning. SubTrack-Grad leverages information from the orthogonal complement of the subspace to enhance subspace estimation through simple linear operations, making it more computationally efficient than GaLore by avoiding periodic SVD on gradient matrices while maintaining the same memory efficiency. Furthermore, its ability to perform simultaneous weight updates provides significant performance improvements over BAdam. Finally, SubTrack-Grad dynamically adapts to the gradient subspace changes, reducing abrupt shifts for faster convergence, in some cases achieving better performance than full-rank finetuning! An overview of SubTrack-Grad is provided in Figure 1, and our main contributions are as follows:
  • We introduce SubTrack-Grad, a memory- and time-efficient full-parameter training that maintains performance by tracking the gradient subspace using Grassmannian geometry.
  • We show that SubTrack-Grad outperforms BAdam, and delivers performance on par with, or superior to GaLore, while significantly reducing runtime with maintaining GaLore’s memory gains.
  • We show that tracking the gradient subspace helps reduce abrupt changes in the optimization process, thereby accelerating convergence.
  • We prove that our method aligns with GaLore’s convergence guarantees while enabling more frequent subspace updates by exercising greater control over subspace adjustments through the use of prior knowledge and estimation errors.
2 Related Works
Several works aim to improve the efficiency of training and fine-tuning LLMs, addressing a growing demand as their popularity rapidly increases. LoRA (Hu et al., 2021), a widely recognized method for reducing the number of trainable parameters, that projects model weights into a lower-dimensional space, resulting in two trainable low-rank matrices. This approach significantly reduces memory requirements for fine-tuning LLMs. Dettmers et al. (2024) build on LoRA by employing quantization techniques and paged optimizers to further reduce memory usage. Additionally, Yaras et al. (2024) introduce Deep LoRA, which uses deep matrix factorization for low-rank optimization, addressing overfitting issues and reducing the need for precise tuning of the rank parameter. Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024). Miles et al. (2024) propose compressing intermediate activation vectors and reconstructing them during backpropagation to enhance memory efficiency. Additionally, Hao et al. (2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient. BAdam (Luo et al., 2024) leverages the block coordinate descent (BCD) framework to reduce memory consumption while maintaining capabilities comparable to Adam.
Several approaches aim to reduce memory consumption in optimizers, as optimizers like Adam (Kingma & Ba, 2017) account for a significant portion of memory usage due to their storage of element-wise states to improve the optimization process (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Rajabi & Rambhatla, 2024c, a). MicroAdam (Modoranu et al., 2024) tackles this issue by compressing the gradient space for optimization and utilizing the resulting compression error through feedback loops to improve the optimization process. Adam-mini (Zhang et al., 2024) partitions model parameters into blocks, assigning a single learning rate to each block. This design achieves a significant reduction in memory usage while maintaining performance. Gur-Ari et al. (2018) suggests that a substantial portion of gradients lies within a small, largely consistent subspace, a finding also reported by other studies, including Schneider et al. (2024); Yaras et al. (2023). GaLore (Zhao et al., 2024) leverages this property to reduce the memory requirements of optimizers by projecting gradients into a lower-dimensional subspace and then projecting them back for full parameter tuning. This approach has been integrated with other methods to further reduce memory usage during training and fine-tuning of LLMs (Li et al., 2024; Rajabi & Rambhatla, 2024b). However, not all layers’ gradients evolve within a stable low-rank subspace. Jaiswal et al. (2024) identifies layers with constantly changing gradients where low-rank projection may be inefficient for tuning. By analyzing the distribution of singular values across different layers, they select those that evolve within a small subspace for fine-tuning while freezing the remaining layers. Gradient Structured Sparsification (Grass) (Muhamed et al., 2024) further reduces memory usage by applying sparse projection matrices to the gradient, transforming the gradient matrix into a sparse space. This approach leverages sparse representations to significantly decrease the memory footprint. In Ramesh et al. (2024) the authors propose an approach that dynamically selects and updates a small subset of parameters, for a faster and more memory-efficient training without altering the model’s structure.
A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve. Balzano et al. (2011) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed. Zhang & Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces. Furthermore, Blocker et al. (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively. Also, Rajabi & Rambhatla (2024d) have applied Grassmannian-based subspace tracking on gradient matrices to increase the efficiency of training.
3 SubTrack-Grad: Tracking the Gradient Subspace
Since gradients typically evolve within a small subspace, compressing this space can significantly reduce the memory footprint of the optimizer. As demonstrated in Zhao et al. (2024) whenever the gradient takes the general form
$\small G=\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i}$ (1)
where $i$ denotes the batch index, and $B_{i}$ and $C_{i}$ are positive semi-definite (PSD) matrices, this gradient can be projected onto a small subspace that remains nearly stable while ensuring that the optimization process continues to converge, as discussed in Section 4 and Appendix A. However, the subspace of the gradient is not always stable, making it crucial to track its changes. GaLore (Zhao et al., 2024) addresses this by periodically performing SVD on gradient matrices, while keeping the update frequency low to align with the assumption of a stable subspace, which is used in its convergence guarantee. This approach poses several challenges: 1) not all gradients converge to a stable subspace within a few iterations, 2) SVD is computationally expensive, and increasing the frequency of updates to capture changes significantly raises both run-time and environmental costs, and 3) increasing the update frequency contradicts the assumption of a stable subspace, as SVD is sensitive to noise and does not account for the previously computed subspace or estimation error to regulate the extent of change applied.
We propose SubTrack-Grad, a memory- and time-efficient method that allows full and simultaneous parameter tuning. SubTrack-Grad utilizes the estimation error and the previously computed subspace to adjust the core subspace. It effectively controls the amount change, allowing more frequent updates without compromising the stability assumption; while, as illustrated in Figure 4, increasing the update frequency will not drastically increase the runtime. The subspace is initialized using SVD as follows:
$\small\begin{split}G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top},\\ P_{0}=[u_{1},u_{2},...,u_{r}],\quad Q_{0}=[v_{1},v_{2},...,v_{r}].\end{split}$ (2)
Here, $G_{0}$ is an $m\times n$ gradient matrix at step 00, $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank. In each optimization step, the gradients are projected onto the subspace of the left singular vectors if $m\leq n$, or onto the right singular vectors otherwise, thereby optimizing memory usage (Zhao et al., 2024). After the optimizer processes the low-rank gradient, it gets projected back for a normal weight update. From now on, we assume $m\leq n$ without loss of generality, implying that $S_{0}=P_{0}$, an $m\times r$ orthonormal matrix whose columns span the underlying subspace.
At each step $t$, the matrix $S_{t}$ represents the subspace at that step, and projects the gradient matrix $G_{t}$ onto that subspace by $\widetilde{G}_{t}=S_{t}^{\top}G_{t}$; where $\widetilde{G}_{t}$, the projection of the gradient onto a rank-$r$ subspace, will be an $r\times n$ matrix. The optimizer then performs within this low-rank space, which substantially reduces the number of state parameters and thus the memory footprint. The optimizer outputs $\widetilde{G}_{t}^{O}$, which is then projected back to the original space by $\widehat{G}_{t}=S_{t}\rho_{t}\widetilde{G}_{t}^{O}$, where $\rho_{t}$ represents the entry-wise regularizer used in the optimizer, to be passed to the network.
Algorithm 1 SubTrack-Grad
0:  Sequence of $m\times n$ gradients $G_{t}$ with $m\leq n$ (w.l.o.g.), step-size $\eta$, rank $r$, subspace update steps k
  Initialize Subspace via SVD Decomposition: $P_{0}\leftarrow U[:,:r]$ , where $U,S,V\leftarrow\text{SVD}(G_{0})$ $S_{0}\leftarrow P_{0}$
  for $t=1,\ldots,T$ do
     if $t\mod k=0$ then
       Update subspace: $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$ $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
     else
       Keep using previous subspace: $S_{t}=S_{t-1}$
     end if
     Return final projected gradient to the optimizer: $S_{t}^{\top}G_{t}$
  end for
As previously discussed, the gradient does not always evolve within a stable subspace; hence, $S_{t}$, the orthonormal matrix spanning the core subspace, must be appropriately updated. SubTrack-Grad updates this subspace by moving along Grassmannian geodesics, and leverages the previously computed subspace and the current estimation error to minimize abrupt changes and noise effects, as illustrated in Figure 3.
We frame the problem of identifying the subspace as selecting the appropriate element from the Grassmannian, the set of all $d$-dimensional subspaces within an $n$-dimensional vector space (Bendokat et al., 2024). The subspace is updated after $k$ steps, the subspace update interval, and our objective is to minimize the Euclidean distance between the subspace and the current observed gradient in an update step. The corresponding cost function for this aim is defined as:
$\small F(S_{t})=\min_{A}\|S_{t}A-G_{t}\|^{2}_{F},$ (3)
where $A$ is the solution to the least squares problem. The derivative of (3) with respect to $S_{t}$ is given in (4), and the residual $R={G_{t}}-S_{t}A$ lies in the orthogonal complement of $S_{t}$. To update the subspace in the appropriate direction, we calculate the tangent vector $\nabla F$ on the Grassmannian, as shown in (5) based on Edelman et al. (1998), where the second equality holds because $R$ is orthogonal to $S_{t}S_{t}^{\top}$.
$\small\frac{\partial F}{\partial S_{t}}=2(S_{t}A-G_{t})A^{\top}=-2RA^{\top}$ (4)
$\small\nabla F=(I-S_{t}S_{t}^{\top})\frac{\partial F}{\partial S_{t}}=\frac{\partial F}{\partial S_{t}}=-2RA^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ (5)
The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement. However, to minimize changes to the subspace, SubTrack-Grad first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (5). This approximation is then used to update the subspace.
As shown by Edelman et al. (1998); Bendokat et al. (2024), from which we have included the necessary definitions and theorem in section 4, we can move along a Grassmannian geodesic guided by the computed tangent vector, taking a step of size $\eta$, as presented in (6).
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$ (6)
This update rule preserves the orthonormality of $S_{t+1}$, ensuring it remains on the Grassmannian. The last term in (5), is required when using thin-SVD; it projects the previous subspace onto the orthogonal complement of $\widehat{V}_{F}$, ensuring that the portion of $S_{t}$ which has not been updated in this step is still included. By leveraging the geometry of the Grassmannian, SubTrack-Grad effectively tracks the underlying subspace of the gradient space, and the pseudo-code for this method is provided in Algorithm 1.
4 Theoretical Analysis
In this section, we analyze the convergence of SubTrack-Grad using theoretical analysis. To begin, the weights update rule of SubTrack-Grad is as follows:
$\small W_{t}=W_{0}+\sum_{t^{\prime}=0}^{t^{\prime}=t-1}\widehat{G}_{t^{\prime}}$ (7)
As previously mentioned, we use left projection if $m\leq n$, where $m$ and $n$ are the dimensions of the gradient matrix, and vice versa. Thus, $\widehat{G}_{t^{\prime}}$ can be computed as shown in (8).
$\small\widehat{G}_{t^{\prime}}=\begin{cases}S_{t^{\prime}}\rho_{t^{\prime}}(S_{t^{\prime}}^{\top}G_{t^{\prime}}),&\text{if $m\leq n$}\\ \rho_{t^{\prime}}(G_{t^{\prime}}S_{t^{\prime}})S_{t^{\prime}}^{\top},&\text{otherwise}\end{cases}$ (8)
Here, $S_{t^{\prime}}$ is the projection matrix that projects the gradient onto the subspace, and $\rho_{t^{\prime}}$ is representing the entry-wise regularizer used in the optimizer. If we use the full projection, then $\widehat{G}_{t^{\prime}}$ will be computed as shown in (9); where $S_{t^{\prime}}^{l}$ and $S_{t^{\prime}}^{r}$ are the rank-$r$ left and right projection matrices.
$\small\widehat{G}_{t^{\prime}}=S_{t^{\prime}}^{l}\rho_{t^{\prime}}({S_{t^{\prime}}^{l}}^{\top}G_{t^{\prime}}S_{t^{\prime}}^{r}){S_{t^{\prime}}^{r}}^{\top}$ (9)
Definition 4.1 (L-continuity).
A function $f(X)$ has Lipschitz-continuity (L-continuity) if for any $X_{1}$ and $X_{2}$, $\|f(X_{2})-f(X_{1})\|_{F}\leq L\|X_{2}-X_{1}\|_{F}$
{restatable}
[Convergence of SubTrack-Grad]theoremconvergence Suppose gradient has the following form (also (1)) with functions $A_{i}$, $B_{i}$, and $C_{i}$ being L-continuous as per Def. 4.1 with constants $L_{A}$, $L_{B}$, and $L_{C}$ w.r.t. weight matrix $W_{t}$; and $\|W_{t}\|_{F}\leq M$; where $W_{t}$ denotes the weight matrix at step $t$, and $M$ is a scalar value,
$\small G=\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i}.$
Now, define $\widehat{B}_{i,t}=(S_{i,t}^{l})^{\top}B_{i}(W_{t})S_{i,t}^{l}$ and $\widehat{C}_{i,t}=(S_{i,t}^{r})^{\top}C_{i}(W_{t})S_{i,t}^{r}$, where $S_{i,t}^{l}$ and $S_{i,t}^{r}$ are the rank-$r$ left and right projection matrices; $B_{i}(W_{t})$ and $C_{i}(W_{t})$ denote the dependence of $B_{i}$ and $C_{i}$ on the weight matrices $W_{t}$. Further letting $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$, and $\kappa_{t}=\frac{1}{N}\sum_{i}\lambda_{min}(\widehat{B}_{i,t})\lambda_{min}(\widehat{C}_{i,t})$, where $\lambda_{min}(\cdot)$ denotes the minimum eigenvalue over each batch, and $N$ representing the number of samples in a batch. Assuming that the projection matrices remain constant during the training. Then for learning-rate $\mu$ and $min(\kappa_{t})>(L_{A}+2L_{B}L_{C}M^{2})$, the SubTrack-Grad, with $\rho_{t}\equiv 1$ (the element-wise regularizer of the optimizer) satisfies:
$\|P_{t}\|_{F}\leq[1-\mu(\kappa_{t-1}-L_{A}-2L_{B}L_{C}M^{2})]\|P_{t-1}\|_{F}.$
That is, $P_{t}\rightarrow 0$ and SubTrack-Grad converges.
The proof of Theorem 4 is provided in Appendix A, based on Zhao et al. (2024). Note that while both GaLore and SubTrack-Grad assume that the subspace remains unchanged for the proof of convergence, GaLore must limit the number of updates to ensure convergence, as each update can potentially change the entire subspace. In contrast, SubTrack-Grad leverages rank-$1$ updates to the subspace, preventing drastic changes with each update. While a deeper analysis of slowly changing subspaces and their impact on convergence remains an open problem, in practice, this allows SubTrack-Grad to perform more frequent updates.
Here we investigate the Grassmannian update rule presented in (6), which is a direct application of Grassmann geometry (Edelman et al., 1998; Bendokat et al., 2024).
Definition 4.2 (Exponential Map).
The exponential map $\exp_{p}:T_{p}M\to M$ on a Riemannian manifold $M$ is a mapping that assigns the point $\gamma(1)\in M$ to each tangent vector $\Delta\in T_{p}M$, where $T_{p}M$ is the tangent space of $M$ at $p$, and $\gamma$ is the unique geodesic originating at $p$ with initial velocity $\Delta$. This map establishes a relationship between geodesics and the Riemannian exponential, such that $\gamma(t)=\exp_{p}(t\Delta)$ for $t\in\mathbb{R}$.
Definition 4.3 (Stiefel Manifold).
The Stiefel manifold $St(n,p)$, parametrizes the set of all $n\times p$ orthonormal matrices $U$, each representing a rank-$p$ subspace of $\mathbb{R}^{n}$.
Definition 4.4 (Grassmann Manifold).
The Grassmannian manifold $Gr(n,p)$ parametrizes the set of all $p$-dimensional subspaces of $\mathbb{R}^{n}$. Each point can be represented by a projection matrix $P=UU^{T}$, where $U\in St(n,p)$.
{restatable}
[Grassmann Exponential]theoremupdaterule Let $P=UU^{T}\in\text{Gr}(n,p)$ be a point on the Grassmannian, where $U\in\text{St}(n,p)$ is the orthonormal basis of the corresponding subspace. Consider a tangent vector $\Delta\in T_{P}\text{Gr}(n,p)$, and let $\Delta_{U}^{\text{hor}}$ denote the horizontal lift of $\Delta$ to the horizontal space at $U$ in the Stiefel manifold $\text{St}(n,p)$. Suppose the thin SVD of $\Delta_{U}^{\text{hor}}$ is given by $\Delta_{U}^{\text{hor}}=\hat{Q}\Sigma V^{T},$ where $\hat{Q}\in\text{St}(n,r)$, $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the nonzero singular values of $\Delta_{U}^{\text{hor}}$ with $r=\min(p,n-p)$, and $V\in\text{St}(p,r)$. The Grassmann exponential map, representing the geodesic emanating from $P$ in the direction $\Delta$, is given by:
$\small\text{Exp}_{P}^{\text{Gr}}(t\Delta)=\big{[}UV\cos(t\Sigma)V^{T}+\hat{Q}\sin(t\Sigma)V^{T}+UV_{\perp}V_{\perp}^{T}\big{]},$
where $V_{\perp}\in\mathbb{R}^{p\times(p-r)}$ is any orthogonal complement of $V$.
The proof of Theorem 4 can be found in Appendix B. Leveraging this theorem and our notation in section 3, one can easily verify that the subspace update rule is as follows:
$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
This update rule generally converges to a stable subspace if the step size $\eta$ decreases over time (Balzano et al., 2011). However, a decreasing step size can impair the ability to accurately track and adapt to subspace changes. Consequently, SubTrack-Grad uses a constant step size during training and fine-tuning to effectively adjust subspaces. This approach does not hinder convergence, as proved in Theorem 4, which guarantees convergence as long as changes are controlled to maintain the stable subspace assumption.
5 Experiments and Results
Table 1: Evaluating the performance of SubTrack-Grad and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$. All hyperparameters are the same for each rank. The performance is measured via Accuracy for QNLI, MNLI, SST-2, and RTE tasks, F1 for QQP and MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA, after fine-tuning for 30 epochs.
COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg
Full-Rank 62.57 91.03 91.32 77.98 94.27 87.83 92.71 89.21 85.86
BAdam (Luo et al., 2024) 54.44 89.01 91.35 68.59 94.15 87.31 92.49 87.50 83.10
GaLore, r=4 (Zhao et al., 2024) 60.34 90.58 92.58 76.53 94.27 87.12 92.20 87.86 85.18
SubTrack-Grad, r=4 (Ours) 61.07 90.63 92.83 76.89 93.81 87.20 91.73 86.73 85.11
GaLore, r=8 (Zhao et al., 2024) 58.54 90.61 91.30 74.37 94.50 87.34 92.71 87.99 84.67
SubTrack(Last)-Grad, r=8 (Ours) 58.03 90.76 91.81 77.62 94.38 87.15 92.31 87.26 84.91
Table 2: Evaluating the performance of SubTrack-Grad and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with rank 8. All hyperparameters are consistent across different methods. The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, F1 for CB, and Matthews Correlation for AXb, after fine-tuning for 30 epochs.
BoolQ CB COPA WIC WSC AXg Avg
Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30 80.64
GaLore (Zhao et al., 2024) 85.44 88.85 80.00 71.47 63.46 100.00 81.54
BAdam (Luo et al., 2024) 82.51 53.28 59.00 70.38 60.58 51.85 62.93
SubTrack-Grad (Ours) 85.66 88.85 81.00 72.73 63.46 100.00 81.95
Figure 2: Comparing wall-time on different architecture and datasets. (a) shows the memory and wall-time measured on pretraining different Llama-architectures sizing from 60M to 3B. (b) demonstrates wall-time measured on fine-tuning RoBERTa-Large on SuperGLUE tasks. (c) presents wall-time measured on fine-tuning RoBERTa-Base on GLUE tasks. As demonstrated, GaLore can lead to 157% wall-time increase in training Llama with 3B parameters, while SubTrack-Grad significantly reduces this overhead.
To ensure a fair comparison between SubTrack-Grad and baselines, we conducted fine-tuning and pre-training experiments across various models and datasets; to measure performance, wall-time, and memory consumption, while all shared hyperparameters remain consistent.
Fine-Tuning Experiments. RoBERTa-Base and RoBERTa-Large were fine-tuned on GLUE and SuperGLUE tasks respectively; with the results presented in Table 1 and Table 2. Figures 2(b) and 2(c) compare wall-time during fine-tuning, with detailed reports provided in Appendix D, as demonstrated, SubTrack-Grad enjoys the same memory reduction as GaLore. For comparing wall-times, models were fine-tuned up to a number of iterations to exactly include five subspace updates, while the evaluation steps were excluded to provide an accurate comparison. SubTrack-Grad demonstrates significant efficiency improvements compared to GaLore by reducing wall-time by up to 20.57% on GLUE tasks (15% average reduction) and up to 65% on SuperGLUE tasks (22% average reduction). Despite limiting updates to rank-$1$ modifications of the previous subspace, SubTrack-Grad achieves superior or comparable performance compared to GaLore. In addition, BAdam is a computationally efficient method due to optimizing one block at a time; but its inability to update all the parameters simultaneously has caused performance loss, and the loss gets greater as the model size increases. Details of fine-tuning hyperparameters are provided in Appendix E.
Pre-Training Experiments. Different Llama-based architectures were pre-trained on C4 datasets for 10K steps, and the results are reported in Table 3. For wall-time comparison, each architecture was trained for 2K iterations, with the subspace update interval set to 200, ensuring exactly 10 subspace updates. Figure 2(a) illustrates wall-times, with detailed reports provided in Appendix D that indicates same memory consumption between GaLore and SubTrack-Grad. Notably, GaLore incurred a substantial 157% increase in wall-time compared to full-rank training for the model of size 3B, whereas SubTrack-Grad exhibited only a 31% increase—representing a 49% reduction in wall-time compared to GaLore. On average, SubTrack-Grad required 27% less training time across these models. Additionally, BAdam demonstrate significant runtime reduction, but its poor performance in pre-training experiments due to its partial optimization, makes it an unsuitable choice for training from scratch. Additional architectural details for the Llama-based models are provided in Appendix F.
Table 3: Evaluation loss comparison for pre-training different Llama-based architectures on C4 dataset after 10K iterations.
60M 130M 350M 1B 3B Avg
r=128 r=256 r=256 r=512 r=512 -
Full-Rank 3.41 3.25 3.40 4.61 4.52 3.84
GaLore 4.02 3.61 3.62 6.53 6.57 4.87
BAdam 7.86 7.08 7.62 7.28 7.12 7.39
SubTrack-Grad 4.13 3.77 3.82 6.53 6.59 4.97
Time and Space Complexity. Table 4 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$. As discussed earlier, GaLore performs SVD on the gradient matrix to estimate the underlying subspace, whereas SubTrack-Grad uses subspace tracking. Comparing the time complexities of these methods, highlights why SubTrack-Grad is significantly more efficient than GaLore. A breakdown of the subspace update time complexity for SubTrack is shown in Appendix C. As shown in Figure 2(a), for pre-training a Llama model of size 3B, GaLore incurs over 157% wall-time overhead, compared to 31% for SubTrack-Grad. This demonstrates a 49% wall-time reduction compared to GaLore. Additionally, the memory required for storing optimizer states in SubTrack-Grad, is equivalent to GaLore’s.
Table 4: Optimizer’s memory requirement in each method along with the time complexity of involved subspace update operations, considering a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$.
Optimizer Memory Subspace Update Time Complexity
Adam $2mn$-
GaLore $mr+2nr$ $O(nm^{2})$
SubTrack-Grad $mr+2nr$ $O(mnr)$
Reducing Optimization Jumps. We argue that GaLore is susceptible to noise, as it relies on the SVD of a single gradient matrix at each subspace update step. Figure 3 compares SubTrack-Grad and GaLore on the Ackley function to illustrate potential jumps and deviations. These jumps prevent GaLore from reaching the global minimum within 100 steps when the scale factor is set to 1. While the scale factor to 3 enables GaLore to reach that minimum, it also results in larger jumps. Figure 5 in Appendix H illustrates that the overall convergence during both pre-training and fine-tuning is similar to other methods when using SubTrack-Grad.
Figure 3: Comparison of the Adam Optimizer Combined with SubTrack-Grad (a, d) and GaLore (b, e) on the Ackley Function for 100 optimization steps and the subspace update interval set to 10. With scale factor 1, GaLore can not reach the to the global minimum of the function due to its abrupt jumps, and for a scale factor of 3, the length of jumps increase. The 3D plots (c, f) provide visualization of the dynamics of the Ackley function and the optimization paths followed by each method in a single plot.
Runtime Consistency. Figure 4 compares the performance and wall-times of GaLore and SubTrack-Grad across subspace update intervals ranging from 50 to 500 while fine-tuning RoBERTa-Base on the COLA task and pre-training a Llama architecture with 60M parameters on the C4 dataset. The subspace update interval denotes the number of iterations between two updates; thus, increasing this interval reduces the update frequency. GaLore’s runtime significantly increases with more frequent subspace updates, whereas SubTrack-Grad maintains minimal overhead. In Table 15 different performance achieved on COLA dataset for different subspace update intervals are compared, illustrating that increasing the update frequency significantly raises GaLore’s runtime, while the performance of the two methods remains very similar.
Figure 4: Comparing runtimes of GaLore and SubTrack-Grad. RoBERTa-Base is fine-tuned for 10 epochs on COLA and Llama-60M is pre-trained for 2500 iterations on C4 dataset. Subspace update intervals range from 50 to 500 in both experiments. Notice that by increasing subspace update interval, the update frequency actually decreases, as it indicates the number of iterations between two subspace update steps. As demonstrated, GaLore’s wall-time can increase drastically when update frequency increases.
6 Discussion and Conclusion
We proposed a time- and memory-efficient method that projects gradients into a lower-dimensional subspace, and preserves the previously computed subspace to incorporate gradient components from the orthogonal complement to perform rank-$1$ updates. This approach reduces the frequency of abrupt transitions between iterations and leverages the available information effectively.
In some cases, the extent of changes in the subspace may require updates of ranks greater than $1$; the pre-training of Llama-based architectures is a clear example of this need. During our experiments, we observed that applying updates as per (6), using the full-rank SVD of the tangent vector from (5), can hinder convergence if the singular values of the tangent vector become very small. Furthermore, simultaneously updating vector spaces associated with different singular values caused convergence issues in some cases. Therefore, we restricted updates to rank-$1$ in this paper, as this approach still enabled SubTrack-Grad to achieve performance comparable to or better than GaLore. In future work, we plan to explore increasing the rank of updates without compromising convergence. Additionally, dynamically selecting the step size could eliminate the need for manual tuning as a hyperparameter which is also left for future works.
Surprisingly, SubTrack-Grad outperforms full-rank fine-tuning on certain tasks, suggesting that low-rank optimization may also serve as a regularizer. Exploring this property further is also an exciting direction for future work.
Appendix A Convergence of SubTrack-Grad
[…]
proof. To demonstrate that SubTrack-Grad converges to the global minimum during training, we begin by deriving the recursive form of the gradients.
Let $\otimes$ denote the Kronecker product. Then, $vec(AXB)=(B^{\top}\otimes A)vec(X)$.
By applying $vec$ to the gradient form given in the theorem, we obtain:
$g_{t}=vec(G_{t})=vec(\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i})=a_{t}-D_{t}w_{t}$ (10)
where $g_{t}:=vec(G_{t})$, $w_{t}:=vec(W_{t})$, $a_{t}:=\frac{1}{N}\sum_{i}vec(A_{i,t})$, and $D_{t}=\frac{1}{N}\sum_{i}C_{i,t}\otimes B_{i,t}$.
As defined in the theorem, let $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$. Its vectorized form can be expressed using the Kronecker product as follows:
$\begin{gathered}p_{t}=vec(P_{t})=vec({S_{t}^{l}}^{\top}G_{t}S_{t}^{r})=({S_{t}^{r}}^{\top}\otimes{S_{t}^{l}}^{\top})vec(G_{t})\\ \quad={(S_{t}^{r}\otimes S_{t}^{l})}^{\top}vec(G_{t})={(S_{t}^{r}\otimes S_{t}^{l})}^{\top}g_{t}\end{gathered}$ (11)
Now recalling $\widehat{G}_{t}$ from (9), it can be written as:
$\widehat{G}_{t}=S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}S_{t}^{r}{S_{t}^{r}}^{\top}$
Thus, its vectorized form will be:
$\begin{gathered}vec(\widehat{G}_{t})=\widehat{g}_{t}=vec(S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}S_{t}^{r}{S_{t}^{r}}^{\top})=vec(S_{t}^{l}P_{t}{S_{t}^{r}}^{\top})\\ \quad=(S_{t}^{r}\otimes S_{t}^{l})vec(P_{t})=(S_{t}^{r}\otimes S_{t}^{l})p_{t}\end{gathered}$ (12)
This is where the constant subspace assumption becomes necessary. To derive the recursive form of $g_{t}$, we assume that the projection matrices remain fixed throughout training, i.e., $S_{t}^{r}=S^{r}$ and $S_{t}^{l}=S^{l}$. Consequently, we can restate equations (11) and (12) as follows:
$p_{t}={(S^{r}\otimes S^{l})}^{\top}g_{t}$ (13)
$\widehat{g}_{t}=(S^{r}\otimes S^{l})p_{t}$ (14)
Then we can write the recursive form of $g_{t}$:
$\begin{gathered}g_{t}=a_{t}-D_{t}w_{t}=(a_{t}-a_{t-1})+(D_{t-1}-D_{t})w_{t}+a_{t-1}-D_{t-1}w_{t}\\ \quad=e_{t}+a_{t-1}-D_{t-1}(w_{t-1}+\mu\widehat{g}_{t-1})=e_{t}+g_{t-1}-\mu D_{t-1}\widehat{g}_{t-1}\end{gathered}$ (15)
where $e_{t}:=(a_{t}-a_{t-1})+(D_{t-1}-D_{t})w_{t}$.
Note that in deriving (15), we utilized the general form of the weight update rule, $w_{t+1}=w_{t}-\mu g_{t}$, which can be rewritten as $w_{t}=w_{t+1}+\mu g_{t}$. By applying this rule along with (10), we arrive at the second equality in (15) as follows:
$\begin{gathered}g_{t}=a_{t}-D_{t}w_{t}=a_{t}-D_{t}w_{t}-g_{t-1}+g_{t-1}\\ \quad=a_{t}-D_{t}w_{t}-a_{t-1}+D_{t-1}w_{t-1}+a_{t-1}-D_{t-1}w_{t-1}\\ \quad=a_{t}-D_{t}w_{t}-a_{t-1}+D_{t-1}(w_{t}+\mu g_{t-1})+a_{t-1}-D_{t-1}(w_{t}+\mu g_{t-1})\\ \quad=a_{t}-D_{t}w_{t}-a_{t-1}+D_{t-1}w_{t}+\mu D_{t-1}g_{t-1}+a_{t-1}-D_{t-1}w_{t}-\mu D_{t-1}g_{t-1}\\ \quad=a_{t}-a_{t-1}+(D_{t-1}-D_{t})w_{t}+a_{t-1}-D_{t-1}\end{gathered}$
To obtain $p_{t}$ from this recursive formulation, we can left-multiply by ${(S^{r}\otimes S^{l})}^{\top}$, as shown in (14):
$\begin{split}p_{t}={(S^{r}\otimes S^{l})}^{\top}e_{t}+{(S^{r}\otimes S^{l})}^{\top}g_{t-1}-\mu{(S^{r}\otimes S^{l})}^{\top}D_{t-1}\widehat{g}_{t-1}\end{split}$ (16)
Now, based on (13) and (14), $p_{t}$ can be written as:
$p_{t}={(S^{r}\otimes S^{l})}^{\top}e_{t}+p_{t-1}-\mu{(S^{r}\otimes S^{l})}^{\top}D_{t-1}{(S^{r}\otimes S^{l})}p_{t-1}$ (17)
Let define:
$\begin{gathered}\widehat{D}_{t}:={(S^{r}\otimes S^{l})}^{\top}D_{t}(S^{r}\otimes S^{l})=\frac{1}{N}\sum_{i}{(S^{r}\otimes S^{l})}^{\top}(C_{i,t}\otimes B_{i,t})(S^{r}\otimes S^{l})\\ \quad=\frac{1}{N}\sum_{i}({S^{r}}^{\top}C_{i,t}S^{r})\otimes({S^{l}}^{\top}B_{i,t}S^{l})\end{gathered}$ (18)
Then we can expand (17) and show that:
$p_{t}=(I-\mu\hat{D}_{t-1})p_{t-1}+(S^{r}\otimes S^{l})^{\top}e_{t}$ (19)
Note that $S^{l}$ and $S^{r}$ are orthonormal matrices. This is ensured because the subspace is initialized using the SVD of $G_{0}$, and the Grassmannian update rule provided in (6) preserves the orthonormality of the subspace matrices throughout training. Since $S^{l}$ and $S^{r}$ are orthonormal, we have ${S^{l}}^{\top}S^{l}=I$ and ${S^{r}}^{\top}S^{r}=I$. Consequently, we can bound the norm of the second term in (19) as follows:
$\|(S^{r}\otimes S^{l})^{\top}e_{t}\|_{2}=\|vec({S^{l}}^{\top}E_{t}S^{r})\|_{2}=\|{S^{l}}^{\top}E_{t}S^{r}\|_{F}\leq\|E_{t}\|_{F}$ (20)
Here $E_{t}$ is the matrix form of $e_{t}$, and as declared before, $e_{t}:=(a_{t}-a_{t-1})+(D_{t-1}-D_{t})w_{t}$, thus:
$E_{t}:=\frac{1}{N}\sum_{i}(A_{i,t}-A_{i,t-1})+\frac{1}{N}\sum_{i}(B_{i,t-1}W_{t}C_{i,t-1}-B_{i,t}W_{t}C_{i,t})$ (21)
Next, we need to find an upper bound for the norm of each term in (21) to establish an upper bound for $\|E_{t}\|_{F}$. Based on the assumptions of the theorem, $A_{i}$, $B_{i}$, and $C_{i}$ exhibit L-Lipschitz continuity with constants $L_{A}$, $L_{B}$, and $L_{C}$, respectively. Additionally, $\|W_{t}\|_{F}$ is bounded by a scalar $M$. We have:
$\|A_{t}-A_{t-1}\|_{F}$$\leq$ $L_{A}\|W_{t}-W_{t-1}\|_{F}=\mu L_{A}\|\tilde{G}_{t-1}\|_{F}\leq\mu L_{A}\|P_{t-1}\|_{F}$ (22)
In the first equality, we apply (7), while the last equality holds due to (14) and the orthonormality of the projection matrices. The subsequent two inequalities can be derived similarly using these equations.
$\begin{gathered}\|(B_{t}-B_{t-1})W_{t}C_{t-1}\|_{F}\leq L_{B}\|W_{t}-W_{t-1}\|_{F}\|W_{t}\|_{F}\|C_{t-1}\|_{F}\\ \quad=\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}\end{gathered}$ (23)
$\begin{gathered}\|B_{t}W_{t}(C_{t-1}-C_{t})\|_{F}\leq L_{C}\|B_{t}\|_{F}\|W_{t}\|_{F}\|W_{t-1}-W_{t}\|_{F}\\ \quad=\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}\end{gathered}$ (24)
We can now derive the bound for $\|E_{t}\|_{F}$ as follows:
$\begin{gathered}\|E_{t}\|_{F}\leq\mu L_{A}\|\tilde{G}_{t-1}\|_{F}\leq\mu L_{A}\|P_{t-1}\|_{F}+\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}+\mu L_{B}L_{C}M^{2}\|P_{t-1}\|_{F}\\ \quad=\mu(L_{A}+2L_{B}L_{C}M^{2})\|P_{t-1}\|_{F}\end{gathered}$ (25)
To calculate the norm bound for the first term in (19), we first need to establish the bounds for $\widehat{D}_{t}$. This involves estimating the minimum eigenvalue of $\widehat{D}_{t}$.
[…]
this actually provides a lower bound for eigenvalues of $\widehat{D}_{t}$, thus:
$\lambda_{\max}(I-\mu\widehat{D}_{t-1})\leq 1-\frac{\mu}{N}\sum_{i}\gamma_{min,i,t-1}$ (27)
considering the definition of $\kappa_{t}$ in the theorem, we can now easily show that:
$\|P_{t}\|_{F}\leq[1-\mu(\kappa_{t-1}-L_{A}-2L_{B}L_{C}M^{2})]\|P_{t-1}\|_{F}.$
and completing the proof.
While SubTrack-Grad utilizes right/left projections to reduce memory consumption, the proof is presented using both projection matrices to ensure generality. Here, we demonstrate how the proof proceeds under the assumption $m\leq n$ (without loss of generality), which allows the use of the left projection matrix.
Using the left projection matrix, the current formulation of $P_{t}$, defined as $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$, simplifies to $P_{t}={S_{t}^{l}}^{\top}G_{t}$. Similarly, $\widehat{G}_{t}=S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}S_{t}^{r}{S_{t}^{r}}^{\top}$ reduces to $\widehat{G}_{t}=S_{t}^{l}{S_{t}^{l}}^{\top}G_{t}$. From this point, the proof continues by substituting $S_{t}^{r}$ with the identity matrix, allowing the derivation of the vectorized forms of $g_{t}$, $\widehat{g}_{t}$, $p_{t}$, and related terms.
The remainder of the proof remains largely unaffected. It can be readily verified that the recursive formulation of $g_{t}$ is unchanged. Although the definition of $P_{t}$ is modified, it continues to satisfy the bounds required for convergence, ensuring that $P_{t}$ converges to 0 when the left projection matrix is used.
Appendix B Grassmann Exponential
[…]
proof. Using Grassmannina mathematics, we know that every $\Delta\in T_{P}Gr(n,p)$ is of the form
$\Delta=Q\begin{pmatrix}0&B^{T}\\ B&0\end{pmatrix}Q^{T}=\left[Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}Q^{T},P\right]$ (28)
Then the lift of $\Delta\in T_{P}Gr(n,p)$ to $Q=\begin{pmatrix}U&U_{\perp}\end{pmatrix}$ can also be calculated explicitly as follows:
$\Delta^{\text{hor}}_{Q}=[\Delta,P]Q=Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}$ (29)
To resume our proof, we need to define the orthogonal group and specifying its tangent space.
Definition B.1 ( Orthogonal Group).
The orthogonal group $O(n)$ is defined as the set of all $n\times n$ matrices $Q$ over $\mathbb{R}$ such that $Q^{T}Q=QQ^{T}=I_{n}$, where $Q^{T}$ is the transpose of $Q$ and $I_{n}$ is the $n\times n$ identity matrix:
$O(n)=\{Q\in\mathbb{R}^{n\times n}\mid Q^{T}Q=I_{n}=QQ^{T}\}.$
Then the tangent space of the orthogonal group $O(n)$ at a point $Q$, denoted $T_{Q}O(n)$, is defined as the set of matrices of the form $Q\Omega$, where $\Omega\in\mathbb{R}^{n\times n}$ is a skew-symmetric matrix, i.e., $\Omega^{T}=-\Omega$:
$T_{Q}O(n)=\{Q\Omega\mid\Omega\in\mathbb{R}^{n\times n},\Omega^{T}=-\Omega\}.$
The geodesic from $Q\in O(n)$ in direction $Q\Omega\in T_{Q}O(n)$ is calculated via
$\text{Exp}_{Q}^{O}(tQ\Omega)=Q\exp_{m}(t\Omega),$ (30)
If $P\in Gr(n,p)$ and $\Delta\in T_{P}Gr(n,p)$ with $\Delta^{\text{hor}}_{Q}=Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}$, the geodesic in the Grassmannian is therefore
$\text{Exp}^{Gr}_{P}(t\Delta)=\pi^{OG}\left(Q\exp_{m}\left(t\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}\right)\right).$ (31)
where $\pi^{OG}$ is the projection from $O(n)$ to $Gr(n,p)$. If the thin SVD of $B$ is given by
$B=U_{\perp}^{T}\Delta^{\text{hor}}_{U}=U_{\perp}^{T}\hat{Q}\Sigma V^{T}$
with $W:=U_{\perp}^{T}\hat{Q}\in St(n-p,r),\Sigma\in\mathbb{R}^{r\times r},V\in St(p,r)$. Let $W_{\perp},V_{\perp}$ be suitable orthogonal completions. Then,
$\exp_{m}\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}=\begin{pmatrix}V&V_{\perp}&0&0\\ 0&0&W&W_{\perp}\end{pmatrix}\begin{pmatrix}\cos(\Sigma)&0&-\sin(\Sigma)&0\\ 0&I_{p-r}&0&0\\ \sin(\Sigma)&0&\cos(\Sigma)&0\\ 0&0&0&I_{n-p-r}\end{pmatrix}\begin{pmatrix}V^{T}&0\\ V_{\perp}^{T}&0\\ 0&W^{T}\\ 0&W_{\perp}^{T}\end{pmatrix},$
which leads to the desired result when inserted into (31). For more mathematical details, you can refer to Edelman et al. (1998), Bendokat et al. (2024), or other useful resources on Grassmann geometry.
Appendix C Time Complexity Analysis
Table 5 presents the time complexity breakdown for the subspace update step in the SubTrack-Grad algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$. As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (3) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix. This operation has a time complexity of $O(mr^{2})$. Computing the residual and the partial derivative with respect to $S_{t}$ requires $O(mrn)$ and $O(mnr)$ time respectively. This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$. The following operation for the partial derivative is $-2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity. The tangent vector computation (5) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$. The rank-1 approximation step uses largest singular value from the SVD of the $m\times r$ tangent vector, and has time complexity of $O(mr^{2})$. Finally, the update rule as shown in (7) which has a time complexity of $O(mr^{2})$. The overall complexity of the algorithm is dominated by the matrix multiplication calculations of time complexity $O(mnr)$. However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack-Grad is still more efficient with respect to time complexity.
Table 5: Time Complexity for SubTrack-Grad Subspace Update
Computation Step Time
Cost function $O(mr^{2})$
Residual $O(mrn)$
Partial derivative $O(mnr)$
Tangent vector $\Delta F$ $O(m^{2}r)$
Rank-$1$ approximation of $\Delta F$ $O(mr^{2})$
Update rule $O(mr^{2})$
Overall $O(mnr)$
Appendix D Memory and Time Comparison
Table 6 presents the wall-times measured to compare the computational efficiency of SubTrack-Grad and GaLore. Additionally, Table 7 shows the peak memory consumption for each architecture during training, using the hyperparameters detailed in 14, on an NVIDIA A100 GPU with 40GB of memory. These values were used to create the bar plots shown in 2(a). For wall-time comparisons, each architecture was trained for 2,000 steps, ensuring exactly 10 subspace updates while excluding evaluation steps. The number of subspace updates was doubled compared to the fine-tuning experiments because larger models and more complex tasks naturally require more frequent updates. This adjustment allowed for more accurate measurements considering this demand.
Table 6: Wall-time comparison for pre-training Llama-based architectures with varying model sizes on the C4 dataset. The experiments consist of 2000 iterations, corresponding to exactly 10 subspace update steps. The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
60M 130M 350M 1B 3B Avg w.r.t FR w.r.t GaLore
r=128 r=256 r=256 r=512 r=512
Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 997.3 - -
BAdam 511.3 779.2 961.6 798.6 1004.1 811.0 -18.68% -
GaLore 547.8 1094.2 1589.0 1729.5 2715.5 1535.2 +53.94% -
SubTrack-Grad 534.9 1061.0 1465.9 1191.1 1385.7 1127.7 +13.08% -26.54%
Table 7: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset. The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
60M 130M 350M 1B 3B Avg w.r.t FR w.r.t GaLore
r=128 r=256 r=256 r=512 r=512
Full-Rank 16.86 25.32 28.67 18.83 34.92 24.92 - -
BAdam 13.34 20.01 21.11 9.70 15.25 15.88 -36.28% -
GaLore 16.89 25.51 27.85 15.24 26.03 22.30 -10.51% -
SubTrack-Grad 16.89 25.52 27.85 15.24 26.16 22.33 -10.39% +0.13%
Table 8 and Table 9 present the wall-time and peak memory consumption, respectively, for fine-tuning RoBERTa-Base on GLUE tasks. These values were used to generate the bar plots shown in 2(c). For wall-time comparisons, the same settings were used, but evaluation steps were excluded, and fine-tuning was limited to number of iterations to ensure exactly 5 subspace update steps during the process.
Table 8: Wall-time comparison when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$ up to a number of iterations to exactly include 5 subspace update steps. The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg w.r.t FR w.r.t GaLore
Full-Rank 105.0 124.0 142.2 260.1 95.0 146.3 161.3 119.0 144.1 - -
BAdam 92.7 103.3 111.4 177.9 85.9 116.0 122.1 99.1 113.5 -21.24% -
GaLore (r=4) 189.2 194.7 198.0 310.7 185.0 205.7 215.1 187.3 210.7 +46.22% -
SubTrack-Grad (r=4) 145.2 155.7 168.3 286.2 141.2 172.6 189.9 151.3 176.3 +22.48% -16.23
GaLore (r=8) 189.9 195.2 195.9 312.3 183.9 205.1 215.7 187.6 210.7 +46.22% -
SubTrack-Grad (r=8) 145.2 156.1 168.4 286.2 143.2 172.0 188.9 152.2 176.6 +22.35 -16.33%
Table 9: Peak memory consumption when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$. The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg w.r.t FR w.r.t GaLore
Full-Rank 4.13 4.24 4.29 8.16 3.88 7.66 9.50 7.80 6.21 - -
BAdam 2.12 2.15 2.30 3.69 1.99 2.13 2.24 2.84 2.43 -60.87% -
GaLore (r=4) 3.64 4.12 3.85 7.27 3.52 6.84 8.75 6.58 5.57 -10.31% -
SubTrack-Grad (r=4) 3.64 4.12 3.86 7.27 3.37 6.73 8.75 6.58 5.54 -10.79% -0.54%
GaLore (r=8) 3.64 4.12 3.86 7.27 3.53 6.84 8.75 6.58 5.57 -10.31% -
SubTrack-Grad (r=8) 3.65 4.12 3.86 7.27 3.38 6.74 8.75 6.58 5.54 -10.79% -0.54%
Table 10 and Table 11 present the wall-time and peak memory consumption for each method when fine-tuning RoBERTa-Large on SuperGLUE tasks. These values were used to generate the bar plots shown in 2(b). For wall-time comparisons, as with GLUE tasks, evaluation steps were omitted to ensure accurate measurements of fine-tuning time, and the number of iterations was adjusted to include exactly 5 subspace updates during fine-tuning.
Table 10: Wall-time comparison when fine-tuning RoBERTa-Large on SuperGLUEup to a number of iterations to exactly include 5 subspace update steps. The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
BoolQ CB COPA WIC WSC AXg Avg w.r.t FR w.r.t GaLore
Full-Rank 1406.3 171.5 35.1 250.4 216.4 52.1 355.3 - -
BAdam 794.9 93.5 45.9 163.9 137.2 38.1 212.25 -40.26% -
GaLore (r=8) 1536.8 272.2 235.3 483.0 393.2 219.9 523.4 +47.31% -
SubTrack-Grad (r=8) 1468.0 198.1 82.2 326.1 266.7 86.6 404.62 +13.88% -22.69%
Table 11: Peak memory consumption comparison when fine-tuning RoBERTa-Large on SuperGLUE. The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
BoolQ CB COPA WIC WSC AXg Avg w.r.t FR w.r.t GaLore
Full-Rank 20.16 15.15 7.86 7.96 9.11 7.81 11.37 - -
BAdam 10.80 5.78 3.36 3.26 4.57 3.26 5.17 -54.52% -
GaLore (r=8) 19.08 12.53 5.32 5.94 7.26 5.34 9.24 -18.73% -
SubTrack-Grad (r=8) 19.09 12.40 5.12 5.95 7.06 5.35 9.16 -19.44% -0.87%
Appendix E Fine-Tuning RoBERTa
RoBERTa-Base was fine-tuned on GLUE tasks using the hyperparameters detailed in Table 12, matching those reported in the GaLore (Zhao et al., 2024) for rank-$4$ and rank-$8$ subspaces, with a subspace update interval set at 500 iterations.
Table 12: Hyperparameters of fine-tuning RoBERTa-Base on GLUE tasks.
MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Shared Parameters Batch Size 16 16 16 32 16 16 16 16
Learning Rate 1E-05 1E-05 3E-05 3E-05 1E-05 1E-05 1E-05 1E-05
# Epochs 30
Max Seq. Len. 512
SubTrack-Grad SubTrack-Grad Step-Size 0.1 0.001 5.0 5.0 0.01 1.0 8.0 10.0
& GaLore, r = 4 Subspace Update Interval 500
Parameters $\alpha$ 4
SubTrack-Grad SubTrack-Grad Step-Size 0.1 0.1 5.0 13.0 0.1 1.0 5.0 10.0
& GaLore, r = 8 Subspace Update Interval 500
Parameters $\alpha$ 2
BAdam Parameters Block Switch Interval 100
Switch Mode Random
We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al. (2024), as detailed in Table 13, with the exception that we fine-tuned each task for 30 epochs.
Table 13: Hyperparameters of fine-tuning RoBERTa-Large on SuperGLUE tasks.
BoolQ CB COPA WIC WSC AXg
Shared Parameters Batch Size 16
# Epochs 30
Learning Rate 1e-5
Max Seq. Len. 512
SubTrack-Grad & SubTrack-Grad Step-Size 0.1 1.0 50.0 50.0 1 1.0
GaLore Parameters Subspace Update Interval 500 100 100 500 250 100
Rank Config. 8
$\alpha$ 4
BAdam Parameters Block Switch Interval 100 50 50 100 50 50
Switch Mode Random
Appendix F Pre-Training LLama-Based Architectures
To manage resources, we pre-trained all five Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 14. While larger models typically require more iterations, this setup was sufficient for comparing methods rather than creating perfectly optimized models.
Table 14: Hyperparameters of pre-training Llama-based architectures.
60M 130M 350M 1B 3B
Architectural Hidden 512 768 1024 2048 4096
Parameters Intermediate 1376 2048 2736 5461 11008
Heads 8 12 16 24 32
Layers 8 12 24 32 32
Shared Parameters Learning Rate 1e-3 1e-3 1e-3 1e-5 1e-5
Batch Size 128 128 64 8 8
Iterations 10K
Gradient Accumulation 2
Gradient Clipping 1.0
Warmup Steps 1000
scale 0.25
dtype bfloat16
SubTrack & Rank 128 256 256 512 512
GaLore Parameters Subspace Update Interval 200
SubTrack-Grad Step-Size 10
BAdam Parameters Block Swithch Interval 100
Switch Mode Random
Appendix G Time and Performance Consistency
Table 15 demonstrates that increasing the update frequency significantly increases GaLore’s runtime, while the performance of both methods remains comparable. This underscores SubTrack-Grad’s efficiency in reducing runtime without sacrificing performance. Notably, this advantage becomes even more pronounced for tasks that demand more frequent updates.
Table 15: Comparing changes in performance while decreasing the subspace update interval (increasing the frequency of subspace updates) on fine-tuning RoBERTa-Base on COLA task.
500 450 400 350 300 250 200 150 100 50
GaLore’s Wall-Time 458.79 460.19 471.36 483.04 495.83 520.28 556.14 605.1 715.87 1038.79
GaLore’s Performance 0.5358 0.5285 0.5385 0.5356 0.5309 0.54821 0.5332 0.5387 0.5463 0.5463
SubTrack-Grad’s Wall-Time 408.22 408.44 408.47 409.25 409.17 409.76 410.29 411.77 413.35 417.07
SubTrack-Grad’s Performance 0.5332 0.5382 0.5364 0.5385 0.5354 0.5364 0.5416 0.5385 0.5364 0.5673
Appendix H Convergence vs. Wall-Time
Figure 5 depicts the changes in the training loss function relative to wall-time. The results demonstrate that SubTrack-Grad’s wall-time reduction has minimal impact on the learning process, showcasing a convergence pattern comparable to other methods without introducing significant computational overhead.
Figure 5: The figures present changes in training loss relative to wall-time for pre-training the Llama-3B architecture on the C4 dataset and fine-tuning RoBERTa-Large on the BoolQ dataset. They demonstrate that SubTrack maintains the overall learning process for both pre-training and fine-tuning while avoiding significant computational overhead.

### v2

@@
-SubTrack your Grad: Gradient Subspace Tracking for Memory and Time Efficient Full-Parameter LLM Training
+SubTrack++ : Gradient Subspace Tracking for Scalable LLM Training
 Abstract
-Training Large Language Models (LLMs) demand significant time and computational resources due to their large model sizes and optimizer states.
-To overcome these challenges, recent methods, such as BAdam, employ partial weight updates to enhance time and memory efficiency, though sometimes at the cost of performance.
-Others, like GaLore, focus on maintaining performance while optimizing memory usage through full parameter training, but may incur higher time complexity.
-By leveraging the low-rank structure of the gradient and the Grassmannian geometry, we propose SubTrack-Grad, a subspace tracking-based optimization method that efficiently tracks the evolving gradient subspace by incorporating estimation errors and previously identified subspaces.
-SubTrack-Grad delivers better or on-par results compared to GaLore, while significantly outperforming BAdam, which, despite being time-efficient, compromises performance.
-SubTrack-Grad reduces wall-time by up to 20.57% on GLUE tasks (15% average reduction) and up to 65% on SuperGLUE tasks (22% average reduction) compared to GaLore.
-Notably, for a 3B parameter model, GaLore incurred a substantial 157% increase in wall-time compared to full-rank training, whereas SubTrack-Grad exhibited a 31% increase—representing a 49% reduction in wall-time, while enjoying the same memory reductions as GaLore.
-Machine Learning, ICML
-Figure 1: Visualization of SubTrack-Grad: Between subspace updates, gradients are projected onto a fixed subspace.
+Training large language models (LLMs) is highly resource-intensive due to their massive number of parameters and the overhead of optimizer states.
+While recent work has aimed to reduce memory consumption, such efforts often entail trade-offs among memory efficiency, training time, and model performance.
+Yet, true democratization of LLMs requires simultaneous progress across all three dimensions.
+To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to changes in the optimization subspace.
+Additionally, employing recovery scaling, a technique that restores information lost through low-rank projections, further enhances model performance.
+Our method demonstrates SOTA convergence by exploiting Grassmannian geometry and achieves lowest evaluation loss, outperforming the current SOTA while reducing pre-training wall time by 43% and maintaining the memory footprint on a 1B-parameter Llama model.
+1 Introduction
+Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity.
+However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact and carbon footprint.
+(Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
+Several techniques have been proposed to mitigate memory bottlenecks, such as gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020).
+LoRA (Hu et al., 2021) and other low-rank adaptation methods (Dettmers et al., 2024; Hu et al., 2021; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024) have gained popularity by optimizing a reduced set of parameters.
+Such approaches often assume that the full parameter space is inherently low-rank—an assumption that does not always hold and can lead to suboptimal performance.
+In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of partial parameter tuning and reduced accuracy.
+Figure 1: We compare baselines on pre-training a 1B-parameter model.
+(a) SubTrack++ achieves the lowest evaluation loss across all methods.
+(b) Its peak memory usage is significantly lower than APOLLO and LDAdam, and on par with GaLore and Fira.
+(c) In terms of wall-time, SubTrack++ incurs minimal overhead relative to APOLLO and is markedly faster than GaLore, Fira, and LDAdam.
+Overall, SubTrack++ outperforms all baselines in evaluation loss while matching or exceeding them in memory and runtime efficiency.
+However, memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024).
+To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024; Rajabi and Rambhatla, 2024a).
+Leveraging the fact that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating the rank-$r$ approximation via singular value decomposition (SVD)—an approach also used in more recent method, Fira (Chen et al., 2025).
+While SVD offers optimal low-rank approximation (Robert et al., 2025), performing it repeatedly can pose several challenges.
+First, it is computationally intensive, and alternatives like random projections (as in APOLLO (Zhu et al., 2025) and GoLore (He et al., 2025)) or approximation methods for estimating dominant singular values (as in LDAdam (Robert et al., 2025), Online Subspace Descent (Liang et al., 2024) and Rajabi and Rambhatla (2024b)) match or outperform SVD in practice.
+Moreover, SVD is sensitive to noise (Vaswani et al., 2018; He et al., 2025) and tends to degrade in late training stages when gradients are small, often hindering convergence (He et al., 2025).
+Geometry-based methods have demonstrated strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Rajabi and Rambhatla, 2024c; Blocker et al., 2023).
+In this context, using Grassmannian manifolds for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016).
+Grassmannian is the set of $r$-dimensional subspaces in a d-dimensional space, and the ultimate goal of low-rank optimization is to find the proper subspace in this space.
+This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017), outperforming online or robust Principle Component Analysis (PCA) techniques (Li, 2004; Balzano et al., 2018; Chakraborty et al., 2017).
+Their natural robustness against perturbations and strong theoretical guarantees make them particularly well-suited to tracking the evolving gradient subspaces encountered in LLM training.
+(Balzano, 2012; Balzano et al., 2011).
+Figure 2: Visualization of Grassmannian-based subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace.
 The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error.
 The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
-1 Introduction
-Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity.
-However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact due to a larger carbon footprint.
-(Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
-Consequently, there is a pressing need to develop time- and memory-efficient methods to make LLMs more accessible and reduce their environmental impact, while preseving performance.
-Several techniques have been proposed to address memory challenges, including gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020).
-In this context, several approaches aim to reduce memory usage by optimizing a subset of model parameters or operating within a lower-dimensional space (Dettmers et al., 2024; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024; Hu et al., 2021).
-Notably, the well-known method LoRA (Hu et al., 2021), addresses memory efficiency by decomposing weight matrices into two low-rank trainable matrices, enabling the optimization of network parameters within a smaller subspace and significantly reducing memory usage.
-However, LoRA’s effectiveness relies on the assumption that parameters reside in a low-dimensional space, which does not always hold true.
-On the other hand, recent methods including BAdam (Luo et al., 2024) and the one introduced by Ramesh et al.
-(2024) leverage the block coordinate descent (BCD) framework and partition the model into different blocks to optimize one block at a time.
-These methods can effectively reduce memory footprint while avoiding extra computation; however, while these achieve better memory and time efficiency, they suffer from performance degradation due to partial parameter tuning.
-Memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024).
-To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024).
-Prominantly, GaLore (Zhao et al., 2024) reduces the memory consumption of optimizers by projecting gradient matrices into a low-rank subspace and tracking changes through periodic singular value decomposition (SVD) to obtain a rank-$r$ approximation.
-Unlike weight parameters, there are many works showing that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023); using this fact, GaLore decreases optimizer memory footprint while enabling full and simultaneous parameter updates.
-However, their approach faces several challenges.
-First, SVD is computationally expensive, and if the gradient does not evolve within a nearly constant subspace, GaLore must increase the frequency of SVD operations, significantly increasing the amount of computation.
-This is problematic because not all layers’ gradients converge to a stable subspace early in training (Jaiswal et al., 2024).
-Moreover, applying SVD to a single gradient matrix is susceptible to data noise (Vaswani et al., 2018), and GaLore does not leverage 1) information from the orthogonal space as feedback to adjust the subspace (Modoranu et al., 2024) or 2) previously computed subspaces to incorporate past knowledge, which could help mitigate noise effects and improve convergence.
-To this end, we propose Gradient Subspace Tracking (SubTrack-Grad), a Grassmannian-based, time- and memory-efficient method that adjusts the subspace using rank-$1$ updates for full parameter fine-tuning.
-SubTrack-Grad leverages information from the orthogonal complement of the subspace to enhance subspace estimation through simple linear operations, making it more computationally efficient than GaLore by avoiding periodic SVD on gradient matrices while maintaining the same memory efficiency.
-Furthermore, its ability to perform simultaneous weight updates provides significant performance improvements over BAdam.
-Finally, SubTrack-Grad dynamically adapts to the gradient subspace changes, reducing abrupt shifts for faster convergence, in some cases achieving better performance than full-rank finetuning!
-An overview of SubTrack-Grad is provided in Figure 1, and our main contributions are as follows:
-  • We introduce SubTrack-Grad, a memory- and time-efficient full-parameter training that maintains performance by tracking the gradient subspace using Grassmannian geometry.
-  • We show that SubTrack-Grad outperforms BAdam, and delivers performance on par with, or superior to GaLore, while significantly reducing runtime with maintaining GaLore’s memory gains.
-  • We show that tracking the gradient subspace helps reduce abrupt changes in the optimization process, thereby accelerating convergence.
-  • We prove that our method aligns with GaLore’s convergence guarantees while enabling more frequent subspace updates by exercising greater control over subspace adjustments through the use of prior knowledge and estimation errors.
-2 Related Works
-Several works aim to improve the efficiency of training and fine-tuning LLMs, addressing a growing demand as their popularity rapidly increases.
-LoRA (Hu et al., 2021), a widely recognized method for reducing the number of trainable parameters, that projects model weights into a lower-dimensional space, resulting in two trainable low-rank matrices.
-This approach significantly reduces memory requirements for fine-tuning LLMs.
-Dettmers et al.
-(2024) build on LoRA by employing quantization techniques and paged optimizers to further reduce memory usage.
-Additionally, Yaras et al.
-(2024) introduce Deep LoRA, which uses deep matrix factorization for low-rank optimization, addressing overfitting issues and reducing the need for precise tuning of the rank parameter.
-Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024).
-Miles et al.
-(2024) propose compressing intermediate activation vectors and reconstructing them during backpropagation to enhance memory efficiency.
-Additionally, Hao et al.
-(2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient.
-BAdam (Luo et al., 2024) leverages the block coordinate descent (BCD) framework to reduce memory consumption while maintaining capabilities comparable to Adam.
-Several approaches aim to reduce memory consumption in optimizers, as optimizers like Adam (Kingma & Ba, 2017) account for a significant portion of memory usage due to their storage of element-wise states to improve the optimization process (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Rajabi & Rambhatla, 2024c, a).
-MicroAdam (Modoranu et al., 2024) tackles this issue by compressing the gradient space for optimization and utilizing the resulting compression error through feedback loops to improve the optimization process.
-Adam-mini (Zhang et al., 2024) partitions model parameters into blocks, assigning a single learning rate to each block.
-This design achieves a significant reduction in memory usage while maintaining performance.
-Gur-Ari et al.
-(2018) suggests that a substantial portion of gradients lies within a small, largely consistent subspace, a finding also reported by other studies, including Schneider et al.
-(2024); Yaras et al.
+To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient method for tracking the gradient subspace.
+This way, instead of reconstructing low-rank approximations via expensive SVD or random projections, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection.
+We also incorporate subspace shifts into Adam’s first and second momentum update rules, ensuring proper alignment with coordinate changes using a Projection-Aware Optimizer (Robert et al., 2025).
+Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection to boost performance by utilizing the scaling information of the low-rank optimizer Chen et al.
+(2025); Zhu et al.
 (2025).
-GaLore (Zhao et al., 2024) leverages this property to reduce the memory requirements of optimizers by projecting gradients into a lower-dimensional subspace and then projecting them back for full parameter tuning.
-This approach has been integrated with other methods to further reduce memory usage during training and fine-tuning of LLMs (Li et al., 2024; Rajabi & Rambhatla, 2024b).
-However, not all layers’ gradients evolve within a stable low-rank subspace.
-Jaiswal et al.
-(2024) identifies layers with constantly changing gradients where low-rank projection may be inefficient for tuning.
-By analyzing the distribution of singular values across different layers, they select those that evolve within a small subspace for fine-tuning while freezing the remaining layers.
-Gradient Structured Sparsification (Grass) (Muhamed et al., 2024) further reduces memory usage by applying sparse projection matrices to the gradient, transforming the gradient matrix into a sparse space.
-This approach leverages sparse representations to significantly decrease the memory footprint.
-In Ramesh et al.
-(2024) the authors propose an approach that dynamically selects and updates a small subset of parameters, for a faster and more memory-efficient training without altering the model’s structure.
-A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve.
-Balzano et al.
-(2011) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed.
-Zhang & Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces.
-Furthermore, Blocker et al.
-(2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively.
-Also, Rajabi & Rambhatla (2024d) have applied Grassmannian-based subspace tracking on gradient matrices to increase the efficiency of training.
-3 SubTrack-Grad: Tracking the Gradient Subspace
-Since gradients typically evolve within a small subspace, compressing this space can significantly reduce the memory footprint of the optimizer.
-As demonstrated in Zhao et al.
-(2024) whenever the gradient takes the general form
-$\small G=\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i}$ (1)
-where $i$ denotes the batch index, and $B_{i}$ and $C_{i}$ are positive semi-definite (PSD) matrices, this gradient can be projected onto a small subspace that remains nearly stable while ensuring that the optimization process continues to converge, as discussed in Section 4 and Appendix A.
-However, the subspace of the gradient is not always stable, making it crucial to track its changes.
-GaLore (Zhao et al., 2024) addresses this by periodically performing SVD on gradient matrices, while keeping the update frequency low to align with the assumption of a stable subspace, which is used in its convergence guarantee.
-This approach poses several challenges: 1) not all gradients converge to a stable subspace within a few iterations, 2) SVD is computationally expensive, and increasing the frequency of updates to capture changes significantly raises both run-time and environmental costs, and 3) increasing the update frequency contradicts the assumption of a stable subspace, as SVD is sensitive to noise and does not account for the previously computed subspace or estimation error to regulate the extent of change applied.
-We propose SubTrack-Grad, a memory- and time-efficient method that allows full and simultaneous parameter tuning.
-SubTrack-Grad utilizes the estimation error and the previously computed subspace to adjust the core subspace.
-It effectively controls the amount change, allowing more frequent updates without compromising the stability assumption; while, as illustrated in Figure 4, increasing the update frequency will not drastically increase the runtime.
-The subspace is initialized using SVD as follows:
-$\small\begin{split}G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top},\\ P_{0}=[u_{1},u_{2},...,u_{r}],\quad Q_{0}=[v_{1},v_{2},...,v_{r}].\end{split}$ (2)
-Here, $G_{0}$ is an $m\times n$ gradient matrix at step 00, $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
-In each optimization step, the gradients are projected onto the subspace of the left singular vectors if $m\leq n$, or onto the right singular vectors otherwise, thereby optimizing memory usage (Zhao et al., 2024).
-After the optimizer processes the low-rank gradient, it gets projected back for a normal weight update.
-From now on, we assume $m\leq n$ without loss of generality, implying that $S_{0}=P_{0}$, an $m\times r$ orthonormal matrix whose columns span the underlying subspace.
-At each step $t$, the matrix $S_{t}$ represents the subspace at that step, and projects the gradient matrix $G_{t}$ onto that subspace by $\widetilde{G}_{t}=S_{t}^{\top}G_{t}$; where $\widetilde{G}_{t}$, the projection of the gradient onto a rank-$r$ subspace, will be an $r\times n$ matrix.
-The optimizer then performs within this low-rank space, which substantially reduces the number of state parameters and thus the memory footprint.
-The optimizer outputs $\widetilde{G}_{t}^{O}$, which is then projected back to the original space by $\widehat{G}_{t}=S_{t}\rho_{t}\widetilde{G}_{t}^{O}$, where $\rho_{t}$ represents the entry-wise regularizer used in the optimizer, to be passed to the network.
-Algorithm 1 SubTrack-Grad
-0:  Sequence of $m\times n$ gradients $G_{t}$ with $m\leq n$ (w.l.o.g.), step-size $\eta$, rank $r$, subspace update steps k
-  Initialize Subspace via SVD Decomposition: $P_{0}\leftarrow U[:,:r]$ , where $U,S,V\leftarrow\text{SVD}(G_{0})$ $S_{0}\leftarrow P_{0}$
-  for $t=1,\ldots,T$ do
-     if $t\mod k=0$ then
-       Update subspace: $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$ $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
-     else
-       Keep using previous subspace: $S_{t}=S_{t-1}$
-     end if
-     Return final projected gradient to the optimizer: $S_{t}^{\top}G_{t}$
-  end for
-As previously discussed, the gradient does not always evolve within a stable subspace; hence, $S_{t}$, the orthonormal matrix spanning the core subspace, must be appropriately updated.
-SubTrack-Grad updates this subspace by moving along Grassmannian geodesics, and leverages the previously computed subspace and the current estimation error to minimize abrupt changes and noise effects, as illustrated in Figure 3.
-We frame the problem of identifying the subspace as selecting the appropriate element from the Grassmannian, the set of all $d$-dimensional subspaces within an $n$-dimensional vector space (Bendokat et al., 2024).
-The subspace is updated after $k$ steps, the subspace update interval, and our objective is to minimize the Euclidean distance between the subspace and the current observed gradient in an update step.
-The corresponding cost function for this aim is defined as:
+To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024), FiRA (Chen et al., 2025), LDAdam (Robert et al., 2025)), while maintaining GaLore’s memory footprint.
+It outperforms online PCA methods like Online Subspace Descent (Liang et al., 2024) in subspace tracking and achieves SOTA convergence and evaluation loss across all strong baselines; Comparison provided in Figure 1.
+2 SubTrack++
+Challenges of Low-Rank Optimization.
+Projecting gradients into a low-rank subspace reduces memory overhead and enables scalable LLM training, but it introduces important trade-offs.
+First, gradient subspaces can be unstable and require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive.
+To address this, recent work has explored cheaper approximations and random projections (Robert et al., 2025; Liang et al., 2024; Zhu et al., 2025; He et al., 2025).
+Furthermore, optimizers like Adam assume a fixed coordinate system, so any changes in the subspace must be reflected in their internal states to preserve consistent momentum updates.
+Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can lead to improved performance (Chen et al., 2025; Zhu et al., 2025).
+Overview.
+We propose SubTrack++ , a memory- and time-efficient method for full, simultaneous parameter tuning that embeds geometric insights into low-rank optimization.
+It improves efficiency and performance via three core components: 1) Subspace Tracking—Grassmannian-based updates that refine projections using estimation error and subspace history; 2) Projection-Aware Optimizer—adjusting Adam’s internal states to account for shifting subspaces; and 3) Recovery Scaling—rescaling discarded gradient components to reclaim lost information.
+Subspace Tracking.
+In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian—the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024)—which is the formal definition and goal of low-rank projections.
+This geometric perspective offers three key benefits: 1) By moving along Grassmann geodesics, it refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in (Zhao et al., 2024; Chen et al., 2025; He et al., 2025).
+2) Updates rely on lightweight algebraic operations, minimizing time and memory costs.
+3) Controlled subspace shifts improve robustness against noise and abrupt changes.
+The initial subspace is computed using SVD as shown in (1). $G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 00; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
+$\small G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top}$ (1)
+At each step, gradients are projected onto the subspace of left singular vectors if $m\leq n$, and right singular vectors otherwise; optimizing memory usage (Zhao et al., 2024).
+We assume $m\leq n$ without loss of generality, so the subspace is represented by $S_{t}\in\mathbb{R}^{m\times r}$ ($S_{0}=[u_{1},...,u_{r}]$), an orthonormal basis spanning the top-$r$ directions.
+The gradient is projected as $\widetilde{G}_{t}=S_{t}^{\top}G_{t}\in\mathbb{R}^{r\times n}$, and the optimizer operates in this reduced space, significantly lowering memory and state overhead.
+To track subspace drift, we adjust the core subspace every $k$ steps, defined as the subspace update interval, by minimizing its Euclidean distance to the current gradient.
+The objective is formalized in the cost function (2).
 $\small F(S_{t})=\min_{A}\|S_{t}A-G_{t}\|^{2}_{F},$ (2)
@@
 The derivative of (2) with respect to $S_{t}$ is given in (3), and the residual $R={G_{t}}-S_{t}A$ lies in the orthogonal complement of $S_{t}$.
-To update the subspace in the appropriate direction, we calculate the tangent vector $\nabla F$ on the Grassmannian, as shown in (5) based on Edelman et al.
+To update the subspace, we calculate the tangent vector $\nabla F$ on the Grassmannian, as shown in (4) based on Edelman et al.
 (1998), where the second equality holds because $R$ is orthogonal to $S_{t}S_{t}^{\top}$.
@@
 The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement.
-However, to minimize changes to the subspace, SubTrack-Grad first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (5).
+However, to minimize changes to the subspace, SubTrack++ first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (4).
 This approximation is then used to update the subspace.
@@
 This update rule preserves the orthonormality of $S_{t+1}$, ensuring it remains on the Grassmannian.
-The last term in (5), is required when using thin-SVD; it projects the previous subspace onto the orthogonal complement of $\widehat{V}_{F}$, ensuring that the portion of $S_{t}$ which has not been updated in this step is still included.
-By leveraging the geometry of the Grassmannian, SubTrack-Grad effectively tracks the underlying subspace of the gradient space, and the pseudo-code for this method is provided in Algorithm 1.
+The last term in (4), projects the previous subspace onto the orthogonal complement of $\widehat{V}_{F}$, ensuring that the portion of $S_{t}$ which has not been updated in this step is still included.
+Algorithm 1 SubTrack++
+(Subspace Tracking, Projection-Aware Optimizer, Recovery Scaling, Regular Adam)
+[…]
+  $S_{0}\leftarrow U[:,:r]$ , where $U,S,V\leftarrow\text{SVD}(G_{0})$Initializing First Subspace}
+  for $t=0,\ldots,T$ do
+     if $t$ mod $k==0$ then
+        $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
+        $M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$$\widetilde{G}_{t}=S_{t}^{\top}G_{t}$: low-rank projection of $G_{t}$} $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
+     else
+        $S_{t}=S_{t-1}$
+        $M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot\mathcal{V}_{t-1}+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
+     end if
+      $\widetilde{G}_{t}^{O}=M_{t}\oslash\sqrt{\mathcal{V}_{t}+\epsilon}$, $\widehat{G}_{t}=S_{t}\widetilde{G}_{t}^{O}$$\widetilde{G}_{t}^{O}$: optimizer’s output, $\widehat{G}_{t}$: projected-back gradients }
+[…]
+     if $\frac{\Lambda_{t}}{\Lambda_{t-1}}>\zeta$ then $\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$
+     $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $-\alpha\cdot\Lambda_{t}$
+  end for
+Projection-Aware Optimizer.
+In Adam, the first and second momentum update rule is shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
+$\small M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (6)
+$\small\mathcal{V}_{t}\leftarrow\beta_{2}\cdot\mathcal{V}_{t-1}+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$ (7)
+Inspired by methods such as LDAdam (Robert et al., 2025), we emphasize the importance of updating optimizer states in a projection-aware manner to account for shifting subspaces; otherwise, misaligned projections can distort the optimizer’s performance.
+To address this, at each subspace update step, we modify Adam’s original update rules in (6) and (7), replacing them with projection-aware counterparts represented in (8) and (9), which reflect subspace changes into optimizer statistics (Robert et al., 2025).
+Further details regarding this projections can be found in Appendix C.
+$\small M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (8)
+$\small\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$ (9)
+These projection-aware update rules enables Adam optimizer to track optimization dynamics precisely as the subspace evolves, achieving significantly better practical performance.
+[…]
+$\small W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}-\alpha\cdot\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ (10)
+Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
+$\small\phi_{t}(G_{t})_{i}=\frac{\|\widetilde{G}_{t,:,i}^{O}\|}{\|\widetilde{G}_{t,:,i}\|}$ (11)
+As suggested by Fira (Chen et al., 2025), this additional term must be controlled using a gradient-clipping-like method to ensure training stability.
+Building on their observation, we also limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
+$\small\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$ (12)
+By incorporating a geometry-aware perspective into low-rank optimization and applying this approach across all components of the training pipeline, SubTrack++ demonstrates robust and stable training, and achieves state-of-the-art performance while maintaining minimal memory footprint and wall-time.
+The overall flow of operations in SubTrack++ is illustrated in Algorithm 1.
 3 Theoretical Analysis
-In this section, we analyze the convergence of SubTrack-Grad using theoretical analysis. To begin, the weights update rule of SubTrack-Grad is as follows:
-$\small W_{t}=W_{0}+\sum_{t^{\prime}=0}^{t^{\prime}=t-1}\widehat{G}_{t^{\prime}}$ (7)
+In this section, we analyze the convergence of Grassmannian Subspace Tracking applied in SubTrack++ using theoretical analysis. To begin, the general weights update rule is as follows:
+$\small W_{t}=W_{0}-\alpha\cdot\sum_{t^{\prime}=0}^{t^{\prime}=t-1}\widehat{G}_{t^{\prime}}$ (13)
 As previously mentioned, we use left projection if $m\leq n$, where $m$ and $n$ are the dimensions of the gradient matrix, and vice versa.
@@
 {restatable}
-[Convergence of SubTrack-Grad]theoremconvergence Suppose gradient has the following form (also (1)) with functions $A_{i}$, $B_{i}$, and $C_{i}$ being L-continuous as per Def.
+[Convergence of Grassmannian Subspace Tracking]theoremconvergence Suppose gradient has the following form with functions $A_{i}$, $B_{i}$, and $C_{i}$ being L-continuous as per Def.
 3.1 with constants $L_{A}$, $L_{B}$, and $L_{C}$ w.r.t.
@@
 Assuming that the projection matrices remain constant during the training.
-Then for learning-rate $\mu$ and $min(\kappa_{t})>(L_{A}+2L_{B}L_{C}M^{2})$, the SubTrack-Grad, with $\rho_{t}\equiv 1$ (the element-wise regularizer of the optimizer) satisfies:
+Then for learning-rate $\mu$ and $min(\kappa_{t})>(L_{A}+2L_{B}L_{C}M^{2})$, subspace tracking, with $\rho_{t}\equiv 1$ (the element-wise regularizer of the optimizer) satisfies:
 $\|P_{t}\|_{F}\leq[1-\mu(\kappa_{t-1}-L_{A}-2L_{B}L_{C}M^{2})]\|P_{t-1}\|_{F}.$
-That is, $P_{t}\rightarrow 0$ and SubTrack-Grad converges.
+That is, $P_{t}\rightarrow 0$ and it converges.
 The proof of Theorem 3 is provided in Appendix A, based on Zhao et al.
 (2024).
-Note that while both GaLore and SubTrack-Grad assume that the subspace remains unchanged for the proof of convergence, GaLore must limit the number of updates to ensure convergence, as each update can potentially change the entire subspace.
-In contrast, SubTrack-Grad leverages rank-$1$ updates to the subspace, preventing drastic changes with each update.
-While a deeper analysis of slowly changing subspaces and their impact on convergence remains an open problem, in practice, this allows SubTrack-Grad to perform more frequent updates.
+While both GaLore and SubTrack++ assume the subspace remains unchanged for the proof of convergence, GaLore must limit these updates to ensure convergence, as each update can potentially change the entire subspace.
+In contrast, SubTrack++ leverages rank-$1$ updates to the subspace, preventing drastic changes with each update.
+While a deeper analysis of slowly changing subspaces and their impact on convergence remains an open problem, in practice, this allows SubTrack++ to perform more frequent updates.
 Here we investigate the Grassmannian update rule presented in (5), which is a direct application of Grassmann geometry (Edelman et al., 1998; Bendokat et al., 2024).
@@
 Definition 3.4 (Grassmann Manifold).
-The Grassmannian manifold $Gr(n,p)$ parametrizes the set of all $p$-dimensional subspaces of $\mathbb{R}^{n}$. Each point can be represented by a projection matrix $P=UU^{T}$, where $U\in St(n,p)$.
+The Grassmannian manifold $Gr(n,p)$ parametrizes the set of all $p$-dimensional subspaces of $\mathbb{R}^{n}$. Each point can be represented by a projection matrix $P=UU^{\top}$, where $U\in St(n,p)$.
 {restatable}
-[Grassmann Exponential]theoremupdaterule Let $P=UU^{T}\in\text{Gr}(n,p)$ be a point on the Grassmannian, where $U\in\text{St}(n,p)$ is the orthonormal basis of the corresponding subspace.
+[Grassmann Exponential]theoremupdaterule Let $P=UU^{\top}\in\text{Gr}(n,p)$ be a point on the Grassmannian, where $U\in\text{St}(n,p)$ is the orthonormal basis of the corresponding subspace.
 Consider a tangent vector $\Delta\in T_{P}\text{Gr}(n,p)$, and let $\Delta_{U}^{\text{hor}}$ denote the horizontal lift of $\Delta$ to the horizontal space at $U$ in the Stiefel manifold $\text{St}(n,p)$.
-Suppose the thin SVD of $\Delta_{U}^{\text{hor}}$ is given by $\Delta_{U}^{\text{hor}}=\hat{Q}\Sigma V^{T},$ where $\hat{Q}\in\text{St}(n,r)$, $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the nonzero singular values of $\Delta_{U}^{\text{hor}}$ with $r=\min(p,n-p)$, and $V\in\text{St}(p,r)$.
+Suppose the thin SVD of $\Delta_{U}^{\text{hor}}$ is given by $\Delta_{U}^{\text{hor}}=\hat{Q}\Sigma V^{\top},$ where $\hat{Q}\in\text{St}(n,r)$, $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the nonzero singular values of $\Delta_{U}^{\text{hor}}$ with $r=\min(p,n-p)$, and $V\in\text{St}(p,r)$.
 The Grassmann exponential map, representing the geodesic emanating from $P$ in the direction $\Delta$, is given by:
-$\small\text{Exp}_{P}^{\text{Gr}}(t\Delta)=\big{[}UV\cos(t\Sigma)V^{T}+\hat{Q}\sin(t\Sigma)V^{T}+UV_{\perp}V_{\perp}^{T}\big{]},$
+$\small\text{Exp}_{P}^{\text{Gr}}(t\Delta)=\big{[}UV\cos(t\Sigma)V^{\top}+\hat{Q}\sin(t\Sigma)V^{\top}+UV_{\perp}V_{\perp}^{\top}\big{]},$
 where $V_{\perp}\in\mathbb{R}^{p\times(p-r)}$ is any orthogonal complement of $V$.
@@
 However, a decreasing step size can impair the ability to accurately track and adapt to subspace changes.
-Consequently, SubTrack-Grad uses a constant step size during training and fine-tuning to effectively adjust subspaces.
+Consequently, SubTrack++ uses a constant step size to effectively adjust subspaces.
 This approach does not hinder convergence, as proved in Theorem 3, which guarantees convergence as long as changes are controlled to maintain the stable subspace assumption.
 4 Experiments and Results
-Table 1: Evaluating the performance of SubTrack-Grad and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$.
-All hyperparameters are the same for each rank.
-The performance is measured via Accuracy for QNLI, MNLI, SST-2, and RTE tasks, F1 for QQP and MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA, after fine-tuning for 30 epochs.
-COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg
-Full-Rank 62.57 91.03 91.32 77.98 94.27 87.83 92.71 89.21 85.86
-BAdam (Luo et al., 2024) 54.44 89.01 91.35 68.59 94.15 87.31 92.49 87.50 83.10
-GaLore, r=4 (Zhao et al., 2024) 60.34 90.58 92.58 76.53 94.27 87.12 92.20 87.86 85.18
-SubTrack-Grad, r=4 (Ours) 61.07 90.63 92.83 76.89 93.81 87.20 91.73 86.73 85.11
-GaLore, r=8 (Zhao et al., 2024) 58.54 90.61 91.30 74.37 94.50 87.34 92.71 87.99 84.67
-SubTrack(Last)-Grad, r=8 (Ours) 58.03 90.76 91.81 77.62 94.38 87.15 92.31 87.26 84.91
-Table 2: Evaluating the performance of SubTrack-Grad and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with rank 8.
-All hyperparameters are consistent across different methods.
-The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, F1 for CB, and Matthews Correlation for AXb, after fine-tuning for 30 epochs.
-BoolQ CB COPA WIC WSC AXg Avg
-Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30 80.64
-GaLore (Zhao et al., 2024) 85.44 88.85 80.00 71.47 63.46 100.00 81.54
-BAdam (Luo et al., 2024) 82.51 53.28 59.00 70.38 60.58 51.85 62.93
-SubTrack-Grad (Ours) 85.66 88.85 81.00 72.73 63.46 100.00 81.95
-Figure 2: Comparing wall-time on different architecture and datasets.
-(a) shows the memory and wall-time measured on pretraining different Llama-architectures sizing from 60M to 3B.
-(b) demonstrates wall-time measured on fine-tuning RoBERTa-Large on SuperGLUE tasks.
-(c) presents wall-time measured on fine-tuning RoBERTa-Base on GLUE tasks.
-As demonstrated, GaLore can lead to 157% wall-time increase in training Llama with 3B parameters, while SubTrack-Grad significantly reduces this overhead.
-To ensure a fair comparison between SubTrack-Grad and baselines, we conducted fine-tuning and pre-training experiments across various models and datasets; to measure performance, wall-time, and memory consumption, while all shared hyperparameters remain consistent.
+To ensure fair comparison, we evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
+Pre-Training Experiments.
+We pre-trained several Llama-based models on the C4 dataset, with results in Table 1.
+To ensure a fair comparison, we benchmarked against a diverse set of baselines.
+BAdam (Luo et al., 2024) is notably efficient in both memory and time, but its reliance on partial parameter updates leads to reduced performance.
+GaLore (Zhao et al., 2024) supports full-parameter training with low memory usage, though this comes at the cost of more than 150% increase in wall-time relative to AdamW in large models.
+Table 1: We compare evaluation loss for pre-training Llama-based architectures on the C4 dataset over 10K iterations.
+SubTrack++ outperforms all other baselines in nearly every configuration.
+The best results are marked in bold, with the second-best performance underlined.
+LDAdam could not be run on the 7B configuration due to an out-of-memory error with our available resources.
+60M 130M 350M 1B 3B 7B
+r=128 r=256 r=256 r=512 r=512 r=1024
+Full-Rank 3.41 3.25 3.40 4.61 4.52 4.30
+GaLore (Zhao et al., 2024) 4.02 3.61 3.62 6.53 6.57 5.55
+BAdam (Luo et al., 2024) 7.86 7.08 7.62 7.28 7.12 6.76
+Online Subspace Descent (Liang et al., 2024) 4.18 3.88 4.09 6.79 6.85 5.69
+LDAdam (Robert et al., 2025) 3.52 3.44 3.67 4.70 4.39 OOM
+Fira (Chen et al., 2025) 3.80 3.55 3.56 6.31 6.50 6.83
+SubTrack++ (Ours) 3.43 3.24 3.29 4.51 4.49 4.62
+As another relative baseline, Online Subspace Descent (Liang et al., 2024) uses Online PCA to track gradient subspace.
+LDAdam (Robert et al., 2025) and Fira (Chen et al., 2025) deliver strong performance at the cost of increasing wall-time even more than GaLore.
+Across these methods, SubTrack++ achieves SOTA results while using 14% less memory and running 43% faster than LDAdam, the most accurate baseline on 1B model.
+Hyperparameter details are provided in Appendix G, with a detailed wall-time and memory analysis in Appendix F.
+Figure 3: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ .
+While Grassmannian tracking alone matches GaLore’s step-wise convergence, it significantly reduces wall-time, as shown in Figure LABEL:fig:modular.
+Figure LABEL:fig:modular compares pre-training performance of a 1B-parameter Llama model across several baselines.
+SubTrack++ converges fastest in both steps and wall-time, emphasizing the power of geometry-aware optimization in boosting performance, reducing resource demands, and democratizing access to SOTA training methods.
+Ablation Study.
+We conducted an ablation study to evaluate the individual and combined effects of projection-aware optimizers and recovery scaling, integrated with Grassmannian subspace tracking on the 1B Llama model.
+We use the experimental settings outlined in Table 10.
+As shown in Figure 3, projection-aware optimizer and recovery scaling independently yield substantial improvements over pure Grassmannian subspace tracking, reducing loss from 6.53 to 5.43 and 5.28, respectively; while their combination, SubTrack++ , achieves 4.51 surpassing all baselines.
+Notably, these gains are achieved with minimal increases in runtime and memory.
 Fine-Tuning Experiments.
-RoBERTa-Base and RoBERTa-Large were fine-tuned on GLUE and SuperGLUE tasks respectively; with the results presented in Table 1 and Table 2.
-Figures 2(b) and 2(c) compare wall-time during fine-tuning, with detailed reports provided in Appendix D, as demonstrated, SubTrack-Grad enjoys the same memory reduction as GaLore.
-For comparing wall-times, models were fine-tuned up to a number of iterations to exactly include five subspace updates, while the evaluation steps were excluded to provide an accurate comparison.
-SubTrack-Grad demonstrates significant efficiency improvements compared to GaLore by reducing wall-time by up to 20.57% on GLUE tasks (15% average reduction) and up to 65% on SuperGLUE tasks (22% average reduction).
-Despite limiting updates to rank-$1$ modifications of the previous subspace, SubTrack-Grad achieves superior or comparable performance compared to GaLore.
-In addition, BAdam is a computationally efficient method due to optimizing one block at a time; but its inability to update all the parameters simultaneously has caused performance loss, and the loss gets greater as the model size increases.
-Details of fine-tuning hyperparameters are provided in Appendix E.
-Pre-Training Experiments.
-Different Llama-based architectures were pre-trained on C4 datasets for 10K steps, and the results are reported in Table 3.
-For wall-time comparison, each architecture was trained for 2K iterations, with the subspace update interval set to 200, ensuring exactly 10 subspace updates.
-Figure 2(a) illustrates wall-times, with detailed reports provided in Appendix D that indicates same memory consumption between GaLore and SubTrack-Grad.
-Notably, GaLore incurred a substantial 157% increase in wall-time compared to full-rank training for the model of size 3B, whereas SubTrack-Grad exhibited only a 31% increase—representing a 49% reduction in wall-time compared to GaLore.
-On average, SubTrack-Grad required 27% less training time across these models.
-Additionally, BAdam demonstrate significant runtime reduction, but its poor performance in pre-training experiments due to its partial optimization, makes it an unsuitable choice for training from scratch.
-Additional architectural details for the Llama-based models are provided in Appendix F.
-Table 3: Evaluation loss comparison for pre-training different Llama-based architectures on C4 dataset after 10K iterations.
-60M 130M 350M 1B 3B Avg
-r=128 r=256 r=256 r=512 r=512 -
-Full-Rank 3.41 3.25 3.40 4.61 4.52 3.84
-GaLore 4.02 3.61 3.62 6.53 6.57 4.87
-BAdam 7.86 7.08 7.62 7.28 7.12 7.39
-SubTrack-Grad 4.13 3.77 3.82 6.53 6.59 4.97
+RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE and SuperGLUE tasks; with the results presented in Table 4 and 5, respectively.
+More details and hyperparameters are provided in Appendix E.
 Time and Space Complexity.
 Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$.
-As discussed earlier, GaLore performs SVD on the gradient matrix to estimate the underlying subspace, whereas SubTrack-Grad uses subspace tracking.
-Comparing the time complexities of these methods, highlights why SubTrack-Grad is significantly more efficient than GaLore.
-A breakdown of the subspace update time complexity for SubTrack is shown in Appendix C.
-As shown in Figure 2(a), for pre-training a Llama model of size 3B, GaLore incurs over 157% wall-time overhead, compared to 31% for SubTrack-Grad.
-This demonstrates a 49% wall-time reduction compared to GaLore.
-Additionally, the memory required for storing optimizer states in SubTrack-Grad, is equivalent to GaLore’s.
-Table 4: Optimizer’s memory requirement in each method along with the time complexity of involved subspace update operations, considering a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$.
-Optimizer Memory Subspace Update Time Complexity
+GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration.
+In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira.
+Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods.
+A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D.
+Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.
+Table 2: The optimizer’s state parameter count and subspace update time complexity across baselines, given a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$.
+LDAdam updates the subspace at every iteration, while other methods update it every $k$ steps.
+Optimizer Mem. Subspace Update Time
 Adam $2mn$-
-GaLore $mr+2nr$ $O(nm^{2})$
-SubTrack-Grad $mr+2nr$ $O(mnr)$
-Reducing Optimization Jumps.
-We argue that GaLore is susceptible to noise, as it relies on the SVD of a single gradient matrix at each subspace update step.
-Figure 3 compares SubTrack-Grad and GaLore on the Ackley function to illustrate potential jumps and deviations.
-These jumps prevent GaLore from reaching the global minimum within 100 steps when the scale factor is set to 1.
-While the scale factor to 3 enables GaLore to reach that minimum, it also results in larger jumps.
-Figure 5 in Appendix H illustrates that the overall convergence during both pre-training and fine-tuning is similar to other methods when using SubTrack-Grad.
-Figure 3: Comparison of the Adam Optimizer Combined with SubTrack-Grad (a, d) and GaLore (b, e) on the Ackley Function for 100 optimization steps and the subspace update interval set to 10.
-With scale factor 1, GaLore can not reach the to the global minimum of the function due to its abrupt jumps, and for a scale factor of 3, the length of jumps increase.
-The 3D plots (c, f) provide visualization of the dynamics of the Ackley function and the optimization paths followed by each method in a single plot.
-Runtime Consistency.
-Figure 4 compares the performance and wall-times of GaLore and SubTrack-Grad across subspace update intervals ranging from 50 to 500 while fine-tuning RoBERTa-Base on the COLA task and pre-training a Llama architecture with 60M parameters on the C4 dataset.
-The subspace update interval denotes the number of iterations between two updates; thus, increasing this interval reduces the update frequency.
-GaLore’s runtime significantly increases with more frequent subspace updates, whereas SubTrack-Grad maintains minimal overhead.
-In Table 15 different performance achieved on COLA dataset for different subspace update intervals are compared, illustrating that increasing the update frequency significantly raises GaLore’s runtime, while the performance of the two methods remains very similar.
-Figure 4: Comparing runtimes of GaLore and SubTrack-Grad.
-RoBERTa-Base is fine-tuned for 10 epochs on COLA and Llama-60M is pre-trained for 2500 iterations on C4 dataset.
-Subspace update intervals range from 50 to 500 in both experiments.
-Notice that by increasing subspace update interval, the update frequency actually decreases, as it indicates the number of iterations between two subspace update steps.
-As demonstrated, GaLore’s wall-time can increase drastically when update frequency increases.
+LDAdam $mr+2nr$ $O(mnr)$
+GaLore, Fira $mr+2nr$ $O(nm^{2})$
+SubTrack++ $mr+2nr$ $O(mnr)$
+(a) Ours, SF=1
+(b) GaLore, SF=1
+(c) Ours, SF=3
+(d) GaLore, SF=3
+Figure 5: Comparison of Grassmannian subspace tracking (Ours) (a, c) and GaLore’s SVD (b, d) on the Ackley Function over 100 optimization steps, with a subspace update interval of 10.
+SF stands for scale factor; with a scale factor of 1, GaLore fails to reach the global minimum due to abrupt jumps.
+At a scale factor of 3, while the minimum is reached, the jump length increases.
+This demonstrates SVD’s sensitivity to noise and abrupt changes, highlighting the robustness of our Subspace Tracking method with its controlled subspace updates.
+Robust Subspace Tracking.
+Relying on SVD for subspace updates makes methods sensitive to noise and abrupt changes (He et al., 2025).
+Figure 5 compares Grassmannian subspace tracking with GaLore’s SVD on the Ackley function, highlighting how SVD causes erratic jumps, while our subspace tracking ensures robust optimization.
+GaLore struggles to reach the global minimum within 100 steps at scale factor 1, and although increasing the scale factor to 3 improves performance, it amplifies jumps that hinder convergence in non-convex settings, revealing sensitivity to hyperparameters, noise, and abrupt changes.
+5 Related Works
+Parameter-Efficient Training.
+Several works aim to improve the efficiency of training LLMs, addressing a growing demand as their popularity rapidly increases.
+Popular LoRA (Hu et al., 2021) significantly reduces memory requirements for fine-tuning LLMs by leveraging two low-rank trainable low-rank matrices.
+Dettmers et al.
+(2024) employ quantization techniques and paged optimizers to further reduce memory usage.
+Additionally, Yaras et al.
+(2024) introduce Deep LoRA to address overfitting issues and reducing the need for precise tuning of the rank parameter.
+Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024).
+Miles et al.
+(2024) propose compressing intermediate activations and reconstructing them during backpropagation to enhance memory efficiency.
+Additionally, Hao et al.
+(2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient.
+BAdam (Luo et al., 2024) leverages the block coordinate descent framework to reduce memory consumption while maintaining capabilities comparable to Adam.
+Gradient Low-Rank Projection.
+Several approaches aim to reduce optimizer states, as optimizers like Adam (Kingma and Ba, 2017) account for a significant portion of memory footprint (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022).
+MicroAdam (Modoranu et al., 2024) tackles this by compressing the gradient space and utilizing the compression error through feedback loops.
+Adam-mini (Zhang et al., 2024) partitions model into blocks, assigning a single learning rate to each block to preserve performance while saving memory.
+Gur-Ari et al.
+(2018); Schneider et al.
+(2024); Yaras et al.
+(2023) suggest that a substantial portion of gradients lies within a largely consistent subspace.
+GaLore (Zhao et al., 2024) leverages this fact to reduce the optimizer’s memory by projecting gradients into a low-rank subspace and then projecting them back for full parameter tuning.
+This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024; Rajabi and Rambhatla, 2024b).
+However, not all layers’ gradients evolve within a stable low-rank subspace.
+Jaiswal et al.
+(2024); Rajabi and Rambhatla (2024d) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates.
+Grass (Muhamed et al., 2024) reduces memory usage by applying sparse projection matrices to the gradient.
+Ramesh et al.
+(2024) dynamically select and update a subset of parameters, for a fast and memory-efficient training.
+FIRA (Chen et al., 2025) utilize a norm-based scaling method along with GaLore to maintain performance comparable to full-rank training.
+GoLore (He et al., 2025) addresses GaLore’s convergence issues and employ random projection in latter steps as a solutions.
+LDAdam (Robert et al., 2025) performs optimization within lower-dimensional subspaces, incorporating a projection-aware optimization update rule and a generalized error feedback mechanism.
+APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection.
+Also, Liang et al.
+(2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD, a similar approach is also proposed in Rajabi and Rambhatla (2024a).
+Grassmannian and Subspace Tracking.
+A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve.
+Balzano et al.
+(2011); Rajabi and Rambhatla (2024c) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed.
+Zhang and Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces.
+Furthermore, Blocker et al.
+(2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively.
 6 Discussion and Conclusion
-We proposed a time- and memory-efficient method that projects gradients into a lower-dimensional subspace, and preserves the previously computed subspace to incorporate gradient components from the orthogonal complement to perform rank-$1$ updates.
-This approach reduces the frequency of abrupt transitions between iterations and leverages the available information effectively.
-In some cases, the extent of changes in the subspace may require updates of ranks greater than $1$; the pre-training of Llama-based architectures is a clear example of this need.
-During our experiments, we observed that applying updates as per (6), using the full-rank SVD of the tangent vector from (5), can hinder convergence if the singular values of the tangent vector become very small.
-Furthermore, simultaneously updating vector spaces associated with different singular values caused convergence issues in some cases.
-Therefore, we restricted updates to rank-$1$ in this paper, as this approach still enabled SubTrack-Grad to achieve performance comparable to or better than GaLore.
-In future work, we plan to explore increasing the rank of updates without compromising convergence.
-Additionally, dynamically selecting the step size could eliminate the need for manual tuning as a hyperparameter which is also left for future works.
-Surprisingly, SubTrack-Grad outperforms full-rank fine-tuning on certain tasks, suggesting that low-rank optimization may also serve as a regularizer.
-Exploring this property further is also an exciting direction for future work.
-Appendix A Convergence of SubTrack-Grad
+We propose SubTrack++ , a time- and memory-efficient approach that projects gradients into a low-rank subspace and uses Grassmannian subspace tracking to preserve the computed subspace while incorporating gradient components from the orthogonal complement.
+By integrating projection-aware optimizers that reflect subspace changes in Adam’s internal statistics and utilizing the gradient information lost during low-rank projection, SubTrack++ achieves state-of-the-art convergence and accuracy across all baselines.
+While Grassmannian subspace tracking integrates seamlessly with various optimizers as a plug-and-play module, extending projection-aware optimization beyond the Adam family requires further design and investigation.
+Also benchmarking on models with more than 7B parameters was not feasible regarding our limited time and resources.
+Appendix A Convergence of SubTrack++
 […]
-proof. To demonstrate that SubTrack-Grad converges to the global minimum during training, we begin by deriving the recursive form of the gradients.
+proof. To demonstrate that SubTrack++ converges to the global minimum during training, we begin by deriving the recursive form of the gradients.
 Let $\otimes$ denote the Kronecker product. Then, $vec(AXB)=(B^{\top}\otimes A)vec(X)$.
@@
 and completing the proof.
-While SubTrack-Grad utilizes right/left projections to reduce memory consumption, the proof is presented using both projection matrices to ensure generality.
+While SubTrack++ utilizes right/left projections to reduce memory consumption, the proof is presented using both projection matrices to ensure generality.
 Here, we demonstrate how the proof proceeds under the assumption $m\leq n$ (without loss of generality), which allows the use of the left projection matrix.
@@
 proof. Using Grassmannina mathematics, we know that every $\Delta\in T_{P}Gr(n,p)$ is of the form
-$\Delta=Q\begin{pmatrix}0&B^{T}\\ B&0\end{pmatrix}Q^{T}=\left[Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}Q^{T},P\right]$ (28)
+$\Delta=Q\begin{pmatrix}0&B^{\top}\\ B&0\end{pmatrix}Q^{\top}=\left[Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}Q^{\top},P\right]$ (34)
 Then the lift of $\Delta\in T_{P}Gr(n,p)$ to $Q=\begin{pmatrix}U&U_{\perp}\end{pmatrix}$ can also be calculated explicitly as follows:
-$\Delta^{\text{hor}}_{Q}=[\Delta,P]Q=Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}$ (29)
+$\Delta^{\text{hor}}_{Q}=[\Delta,P]Q=Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}$ (35)
 To resume our proof, we need to define the orthogonal group and specifying its tangent space.
 Definition B.1 (Orthogonal Group).
-The orthogonal group $O(n)$ is defined as the set of all $n\times n$ matrices $Q$ over $\mathbb{R}$ such that $Q^{T}Q=QQ^{T}=I_{n}$, where $Q^{T}$ is the transpose of $Q$ and $I_{n}$ is the $n\times n$ identity matrix:
-$O(n)=\{Q\in\mathbb{R}^{n\times n}\mid Q^{T}Q=I_{n}=QQ^{T}\}.$
-Then the tangent space of the orthogonal group $O(n)$ at a point $Q$, denoted $T_{Q}O(n)$, is defined as the set of matrices of the form $Q\Omega$, where $\Omega\in\mathbb{R}^{n\times n}$ is a skew-symmetric matrix, i.e., $\Omega^{T}=-\Omega$:
-$T_{Q}O(n)=\{Q\Omega\mid\Omega\in\mathbb{R}^{n\times n},\Omega^{T}=-\Omega\}.$
+The orthogonal group $O(n)$ is defined as the set of all $n\times n$ matrices $Q$ over $\mathbb{R}$ such that $Q^{\top}Q=QQ^{\top}=I_{n}$, where $Q^{\top}$ is the transpose of $Q$ and $I_{n}$ is the $n\times n$ identity matrix:
+$O(n)=\{Q\in\mathbb{R}^{n\times n}\mid Q^{\top}Q=I_{n}=QQ^{\top}\}.$
+Then the tangent space of the orthogonal group $O(n)$ at a point $Q$, denoted $T_{Q}O(n)$, is defined as the set of matrices of the form $Q\Omega$, where $\Omega\in\mathbb{R}^{n\times n}$ is a skew-symmetric matrix, i.e., $\Omega^{\top}=-\Omega$:
+$T_{Q}O(n)=\{Q\Omega\mid\Omega\in\mathbb{R}^{n\times n},\Omega^{\top}=-\Omega\}.$
 The geodesic from $Q\in O(n)$ in direction $Q\Omega\in T_{Q}O(n)$ is calculated via
 $\text{Exp}_{Q}^{O}(tQ\Omega)=Q\exp_{m}(t\Omega),$ (36)
-If $P\in Gr(n,p)$ and $\Delta\in T_{P}Gr(n,p)$ with $\Delta^{\text{hor}}_{Q}=Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}$, the geodesic in the Grassmannian is therefore
-$\text{Exp}^{Gr}_{P}(t\Delta)=\pi^{OG}\left(Q\exp_{m}\left(t\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}\right)\right).$ (31)
+If $P\in Gr(n,p)$ and $\Delta\in T_{P}Gr(n,p)$ with $\Delta^{\text{hor}}_{Q}=Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}$, the geodesic in the Grassmannian is therefore
+$\text{Exp}^{Gr}_{P}(t\Delta)=\pi^{OG}\left(Q\exp_{m}\left(t\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}\right)\right).$ (37)
 where $\pi^{OG}$ is the projection from $O(n)$ to $Gr(n,p)$. If the thin SVD of $B$ is given by
-$B=U_{\perp}^{T}\Delta^{\text{hor}}_{U}=U_{\perp}^{T}\hat{Q}\Sigma V^{T}$
-with $W:=U_{\perp}^{T}\hat{Q}\in St(n-p,r),\Sigma\in\mathbb{R}^{r\times r},V\in St(p,r)$. Let $W_{\perp},V_{\perp}$ be suitable orthogonal completions. Then,
-$\exp_{m}\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}=\begin{pmatrix}V&V_{\perp}&0&0\\ 0&0&W&W_{\perp}\end{pmatrix}\begin{pmatrix}\cos(\Sigma)&0&-\sin(\Sigma)&0\\ 0&I_{p-r}&0&0\\ \sin(\Sigma)&0&\cos(\Sigma)&0\\ 0&0&0&I_{n-p-r}\end{pmatrix}\begin{pmatrix}V^{T}&0\\ V_{\perp}^{T}&0\\ 0&W^{T}\\ 0&W_{\perp}^{T}\end{pmatrix},$
+$B=U_{\perp}^{\top}\Delta^{\text{hor}}_{U}=U_{\perp}^{\top}\hat{Q}\Sigma V^{\top}$
+with $W:=U_{\perp}^{\top}\hat{Q}\in St(n-p,r),\Sigma\in\mathbb{R}^{r\times r},V\in St(p,r)$. Let $W_{\perp},V_{\perp}$ be suitable orthogonal completions. Then,
+$\exp_{m}\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}=\begin{pmatrix}V&V_{\perp}&0&0\\ 0&0&W&W_{\perp}\end{pmatrix}\begin{pmatrix}\cos(\Sigma)&0&-\sin(\Sigma)&0\\ 0&I_{p-r}&0&0\\ \sin(\Sigma)&0&\cos(\Sigma)&0\\ 0&0&0&I_{n-p-r}\end{pmatrix}\begin{pmatrix}V^{\top}&0\\ V_{\perp}^{\top}&0\\ 0&W^{\top}\\ 0&W_{\perp}^{\top}\end{pmatrix},$
 which leads to the desired result when inserted into (37).
 For more mathematical details, you can refer to Edelman et al.
-(1998), Bendokat et al.
-(2024), or other useful resources on Grassmann geometry.
-Appendix C Time Complexity Analysis
-Table 5 presents the time complexity breakdown for the subspace update step in the SubTrack-Grad algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$.
+[1998], Bendokat et al.
+[2024], or other useful resources on Grassmann geometry.
+Appendix C Projection-Aware Optimizer
+When projecting into a lower-dimensional space and tracking coordinate changes, we typically use orthonormal projection matrices to represent the subspaces and their multiplications to effect a change of basis.
+While this works well for purely linear operations, Adam’s updates also incorporate non-linear elements.
+Suppose the subspace changes at step $t$, transitioning from the subspace spanned by the orthonormal matrix $S_{t-1}$ to that spanned by $S_{t}$.
+Since both matrices are orthonormal—preserved by the Grassmannian update rule in (5)—the matrix $S_{t}^{\top}S_{t-1}$ represents the change of basis between the two subspaces.
+In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps $t-1$ and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step $t-1$.
+$\mathbf{E}_{t,\beta}\left[.\right]$ denotes the exponential time-weighted expectation at time $t$ with decay rate $\beta$.
+Following the reinterpretation by Robert et al.
+[2025], Adam’s first and second moment estimates can be expressed as $\widetilde{M}_{t}=\mathbf{E}_{t,\beta_{1}}\left[\widetilde{G}_{t}\right]$ and $\widetilde{\mathcal{V}}_{t}=\mathbf{E}_{t,\beta_{2}}\left[(\widetilde{G}_{t})^{2}\right]$, where $\widetilde{G}_{t}$ denotes the low-rank representation of the gradient at time step $t$.
+As shown in (38), the first moment estimate can be transformed under a change of basis using the change-of-basis matrix, $S_{t}^{\top}S_{t-1}$.
+Notably, $\langle\widetilde{G}_{t},e^{i}_{t}\rangle$ gives the $i$th column of $\widetilde{G}_{t}$ when the subspace has the basis $\mathbb{E}_{t}$.
+We use the superscripts to indicate a column of a matrix.
+$\small\mathbf{E}_{t,\beta_{1}}\left[\langle\widetilde{G}_{t},e^{i}_{t}\rangle\right]=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle\mathbf{E}_{t,\beta_{1}}\left[\langle\widetilde{G}_{t},e^{j}_{t-1}\rangle\right]=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle\widetilde{M}_{t}^{j}=\left(S_{t}^{\top}S_{t-1}\widetilde{M}_{t}\right)^{i}$ (38)
+Following the same approach, we can change the basis for the second moment estimate as described in (39).
+$\small\mathbf{E}_{t,\beta_{2}}\left[\left(\langle\widetilde{G}_{t},e^{i}_{t}\rangle\right)^{2}\right]$ $=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\,\mathbf{E}_{t,\beta_{2}}\left[\left(\langle\widetilde{G}_{t},e^{j}_{t-1}\rangle\right)^{2}\right]$ (39)
+$+\sum_{k\neq l}^{r}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\,\mathbf{E}_{t,\beta_{2}}\left[\langle\widetilde{G}_{t},e^{k}_{t-1}\rangle\langle\widetilde{G}_{t},e^{l}_{t-1}\rangle\right]$
+$=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\widetilde{\mathcal{V}}_{t}^{j}$
+$+\sum_{k\neq l}^{r}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}.$
+In transitioning from the first equality to the second, we assume independence among the gradient coordinates.
+This enables us to approximate the covariance using a product of first-order moment estimates.
+This assumption is often reasonable in practice because we compute the SVD of the gradient and maintain an orthonormal subspace projection matrix, updating it along the Grassmannian geodesic to track the optimal subspace.
+Since SVD tends to diagonalize the covariance, the off-diagonal entries are typically negligible.
+Additionally, we clip any negative values to zero to ensure valid (non-negative) variance estimates.
+Moreover, to rewrite the second term in the final equality of (39), we employ the following equation:
+$\sum_{k}\sum_{l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}$ (40)
+$=\sum_{k}\langle e^{i}_{t},e^{k}_{t-1}\rangle^{2}\left(\widetilde{M}_{t}^{k}\right)^{2}\;+\;\sum_{k\neq l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}$
+Given these, we can rewrite (39) as follows:
+$\mathbf{E}_{t,\beta_{2}}\left[\left\langle\widetilde{G}_{t},e_{t}^{i}\right\rangle^{2}\right]$ $=\sum_{j}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\widetilde{\mathcal{V}}_{t}^{j}+$ (41)
+$\left[\sum_{k}\sum_{l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}-\sum_{k}\langle e^{i}_{t},e^{k}_{t-1}\rangle^{2}\left(\widetilde{M}_{t}^{k}\right)^{2}\right]$
+$=\sum_{j}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\left[\widetilde{\mathcal{V}}_{t}^{j}-\left(\widetilde{M}_{t}^{j}\right)^{2}\right]+\left(\langle e^{i}_{t},e^{j}_{t-1}\rangle\widetilde{M}_{t}{}\right)^{2}$
+$=\left(\left(S_{t}^{\top}S_{t-1}\right)^{2}\left[\widetilde{\mathcal{V}}_{t}-\widetilde{M}_{t}^{2}\right]\right)^{i}+\left(\left(S_{t}^{\top}S_{t-1}\widetilde{M}_{t}\right)^{2}\right)^{i}$
+By applying (38) and (41), we can directly derive the update rules for the projection-aware optimizer, as expressed in (8) and (9).
+Appendix D Time Complexity Analysis
+Table 3 presents the time complexity breakdown for the subspace update step in the SubTrack++ algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$.
 As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (2) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix.
@@
 The tangent vector computation (4) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$.
-The rank-1 approximation step uses largest singular value from the SVD of the $m\times r$ tangent vector, and has time complexity of $O(mr^{2})$.
-Finally, the update rule as shown in (7) which has a time complexity of $O(mr^{2})$.
-The overall complexity of the algorithm is dominated by the matrix multiplication calculations of time complexity $O(mnr)$.
-However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack-Grad is still more efficient with respect to time complexity.
-Table 5: Time Complexity for SubTrack-Grad Subspace Update
+Table 3: Time Complexity for SubTrack++ Subspace Update
 Computation Step Time
@@
 Overall $O(mnr)$
-Appendix D Memory and Time Comparison
-Table 6 presents the wall-times measured to compare the computational efficiency of SubTrack-Grad and GaLore.
-Additionally, Table 7 shows the peak memory consumption for each architecture during training, using the hyperparameters detailed in 14, on an NVIDIA A100 GPU with 40GB of memory.
-These values were used to create the bar plots shown in 2(a).
-For wall-time comparisons, each architecture was trained for 2,000 steps, ensuring exactly 10 subspace updates while excluding evaluation steps.
-The number of subspace updates was doubled compared to the fine-tuning experiments because larger models and more complex tasks naturally require more frequent updates.
-This adjustment allowed for more accurate measurements considering this demand.
-Table 6: Wall-time comparison for pre-training Llama-based architectures with varying model sizes on the C4 dataset.
-The experiments consist of 2000 iterations, corresponding to exactly 10 subspace update steps.
-The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
-60M 130M 350M 1B 3B Avg w.r.t FR w.r.t GaLore
-r=128 r=256 r=256 r=512 r=512
-Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 997.3 - -
-BAdam 511.3 779.2 961.6 798.6 1004.1 811.0 -18.68% -
-GaLore 547.8 1094.2 1589.0 1729.5 2715.5 1535.2 +53.94% -
-SubTrack-Grad 534.9 1061.0 1465.9 1191.1 1385.7 1127.7 +13.08% -26.54%
-Table 7: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset.
-The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
-60M 130M 350M 1B 3B Avg w.r.t FR w.r.t GaLore
-r=128 r=256 r=256 r=512 r=512
-Full-Rank 16.86 25.32 28.67 18.83 34.92 24.92 - -
-BAdam 13.34 20.01 21.11 9.70 15.25 15.88 -36.28% -
-GaLore 16.89 25.51 27.85 15.24 26.03 22.30 -10.51% -
-SubTrack-Grad 16.89 25.52 27.85 15.24 26.16 22.33 -10.39% +0.13%
-Table 8 and Table 9 present the wall-time and peak memory consumption, respectively, for fine-tuning RoBERTa-Base on GLUE tasks.
-These values were used to generate the bar plots shown in 2(c).
-For wall-time comparisons, the same settings were used, but evaluation steps were excluded, and fine-tuning was limited to number of iterations to ensure exactly 5 subspace update steps during the process.
-Table 8: Wall-time comparison when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$ up to a number of iterations to exactly include 5 subspace update steps.
-The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
-COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg w.r.t FR w.r.t GaLore
-Full-Rank 105.0 124.0 142.2 260.1 95.0 146.3 161.3 119.0 144.1 - -
-BAdam 92.7 103.3 111.4 177.9 85.9 116.0 122.1 99.1 113.5 -21.24% -
-GaLore (r=4) 189.2 194.7 198.0 310.7 185.0 205.7 215.1 187.3 210.7 +46.22% -
-SubTrack-Grad (r=4) 145.2 155.7 168.3 286.2 141.2 172.6 189.9 151.3 176.3 +22.48% -16.23
-GaLore (r=8) 189.9 195.2 195.9 312.3 183.9 205.1 215.7 187.6 210.7 +46.22% -
-SubTrack-Grad (r=8) 145.2 156.1 168.4 286.2 143.2 172.0 188.9 152.2 176.6 +22.35 -16.33%
-Table 9: Peak memory consumption when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$.
-The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
-COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg w.r.t FR w.r.t GaLore
-Full-Rank 4.13 4.24 4.29 8.16 3.88 7.66 9.50 7.80 6.21 - -
-BAdam 2.12 2.15 2.30 3.69 1.99 2.13 2.24 2.84 2.43 -60.87% -
-GaLore (r=4) 3.64 4.12 3.85 7.27 3.52 6.84 8.75 6.58 5.57 -10.31% -
-SubTrack-Grad (r=4) 3.64 4.12 3.86 7.27 3.37 6.73 8.75 6.58 5.54 -10.79% -0.54%
-GaLore (r=8) 3.64 4.12 3.86 7.27 3.53 6.84 8.75 6.58 5.57 -10.31% -
-SubTrack-Grad (r=8) 3.65 4.12 3.86 7.27 3.38 6.74 8.75 6.58 5.54 -10.79% -0.54%
-Table 10 and Table 11 present the wall-time and peak memory consumption for each method when fine-tuning RoBERTa-Large on SuperGLUE tasks.
-These values were used to generate the bar plots shown in 2(b).
-For wall-time comparisons, as with GLUE tasks, evaluation steps were omitted to ensure accurate measurements of fine-tuning time, and the number of iterations was adjusted to include exactly 5 subspace updates during fine-tuning.
-Table 10: Wall-time comparison when fine-tuning RoBERTa-Large on SuperGLUEup to a number of iterations to exactly include 5 subspace update steps.
-The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
-BoolQ CB COPA WIC WSC AXg Avg w.r.t FR w.r.t GaLore
-Full-Rank 1406.3 171.5 35.1 250.4 216.4 52.1 355.3 - -
-BAdam 794.9 93.5 45.9 163.9 137.2 38.1 212.25 -40.26% -
-GaLore (r=8) 1536.8 272.2 235.3 483.0 393.2 219.9 523.4 +47.31% -
-SubTrack-Grad (r=8) 1468.0 198.1 82.2 326.1 266.7 86.6 404.62 +13.88% -22.69%
-Table 11: Peak memory consumption comparison when fine-tuning RoBERTa-Large on SuperGLUE.
-The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
-BoolQ CB COPA WIC WSC AXg Avg w.r.t FR w.r.t GaLore
-Full-Rank 20.16 15.15 7.86 7.96 9.11 7.81 11.37 - -
-BAdam 10.80 5.78 3.36 3.26 4.57 3.26 5.17 -54.52% -
-GaLore (r=8) 19.08 12.53 5.32 5.94 7.26 5.34 9.24 -18.73% -
-SubTrack-Grad (r=8) 19.09 12.40 5.12 5.95 7.06 5.35 9.16 -19.44% -0.87%
-Appendix E Fine-Tuning RoBERTa
-RoBERTa-Base was fine-tuned on GLUE tasks using the hyperparameters detailed in Table 12, matching those reported in the GaLore (Zhao et al., 2024) for rank-$4$ and rank-$8$ subspaces, with a subspace update interval set at 500 iterations.
+The rank-1 approximation step uses largest singular value from the SVD of the $m\times r$ tangent vector, and has time complexity of $O(mr^{2})$.
+Finally, the update rule as shown in (13) which has a time complexity of $O(mr^{2})$.
+The overall complexity of the algorithm is dominated by the matrix multiplication calculations of time complexity $O(mnr)$.
+However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack++ is still more efficient with respect to time complexity.
+Appendix E Fine-Tuning Experiments
+As described, we examined SubTrack++ on fine-tuning RoBERTa-base and RoBERTa-large models to evaluate them on GLUE and SuperGLUE benchmarks.
+The results for GLUE task is summarized in Table 4, and SuperGLUE in 5.
+Table 4: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for $r=8$.
+The performance is measured via Accuracy for SST-2 and RTE tasks, F1 for MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA.
+The best results are marked in bold, with the second-best performance underlined.
+COLA STS-B MRPC RTE SST-2
+Full-Rank 62.57 91.03 91.32 77.98 94.27
+BAdam [Luo et al., 2024] 54.44 89.01 91.35 68.59 94.15
+GaLore [Zhao et al., 2024] 58.54 90.61 91.30 74.37 94.50
+LDAdam [Robert et al., 2025] 58.81 90.90 92.22 76.53 94.27
+SubTrack++ (Ours) 58.29 90.90 91.84 76.53 90.83
+Table 5: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with $r=8$.
+The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, and F1 for CB.
+The best results are marked in bold, with the second-best performance underlined.
+BoolQ CB COPA WIC WSC AXg
+Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30
+GaLore [Zhao et al., 2024] 85.44 88.85 80.00 71.47 63.46 100.00
+BAdam [Luo et al., 2024] 82.51 53.28 59.00 70.38 60.58 51.85
+LDAdam Robert et al. [2025] 85.75 56.16 80.00 71.00 64.42 70.37
+SubTrack++ (Ours) 84.68 89.66 88.00 70.06 63.46 100.00
 Table 6: Hyperparameters of fine-tuning RoBERTa-Base on GLUE tasks.
-MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
-Shared Parameters Batch Size 16 16 16 32 16 16 16 16
-Learning Rate 1E-05 1E-05 3E-05 3E-05 1E-05 1E-05 1E-05 1E-05
+SST-2 MRPC CoLA RTE STS-B
+Shared Parameters Batch Size 16 16 32 16 16
 # Epochs 30
 Max Seq. Len. 512
-SubTrack-Grad SubTrack-Grad Step-Size 0.1 0.001 5.0 5.0 0.01 1.0 8.0 10.0
-& GaLore, r = 4 Subspace Update Interval 500
-Parameters $\alpha$ 4
-SubTrack-Grad SubTrack-Grad Step-Size 0.1 0.1 5.0 13.0 0.1 1.0 5.0 10.0
-& GaLore, r = 8 Subspace Update Interval 500
-Parameters $\alpha$ 2
-BAdam Parameters Block Switch Interval 100
+Low-Rank Optimizer Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
+Methods Parameters SubTrack Step-Size 0.01 3.0 5.0 15.0 10.0
+Subspace Update Interval 500
+Rank Config 8
+$\alpha$ 2
+BAdam Parameters Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
+Block Switch Interval 100
 Switch Mode Random
-We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al. (2024), as detailed in Table 13, with the exception that we fine-tuned each task for 30 epochs.
 Table 7: Hyperparameters of fine-tuning RoBERTa-Large on SuperGLUE tasks.
@@
 Max Seq. Len. 512
-SubTrack-Grad & SubTrack-Grad Step-Size 0.1 1.0 50.0 50.0 1 1.0
-GaLore Parameters Subspace Update Interval 500 100 100 500 250 100
+Low-Rank Optimizer SubTrack++ Step-Size 0.1 1.0 50.0 100.0 1.0 1.0
+Methods Parameters Subspace Update Interval 500 100 100 500 250 100
 Rank Config. 8
@@
 Switch Mode Random
-Appendix F Pre-Training LLama-Based Architectures
-To manage resources, we pre-trained all five Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 14.
-While larger models typically require more iterations, this setup was sufficient for comparing methods rather than creating perfectly optimized models.
+The hyperparameters for fine-tuning RoBERTa-base are detailed in Table 6, matching those reported in the GaLore [Zhao et al., 2024] for rank-$8$ subspaces, with a subspace update interval set at 500 iterations.
+We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al.
+[2024], as detailed in Table 7, with the exception that we fine-tuned each task for 30 epochs.
+Appendix F Memory and Time Comparison
+Table 8 presents the the peak memory consumption measured to compare SubTrack++ and other baselines.
+It shows that SubTrack++ requires on-par or better memory compared to GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025] while getting state-of-the-art results.
+As detailed in Table 2, all baselines except BAdam [Luo et al., 2024] use the same number of optimizer parameters; therefore, any differences in their peak memory consumption stem from variations in their runtime parameters.
+Table 8: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset.
+The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with .
+SubTrack++ demonstrates better or on-par memory compared to other low-rank methods that allows full-parameter training.
+60M 130M 350M 1B 3B 7B
+r=128 r=256 r=256 r=512 r=512 r=1024
+Full-Rank 16.86 25.32 28.67 18.83 34.92 50.50
+BAdam [Luo et al., 2024] 13.34 20.01 16.45 9.18 14.75 22.35
+GaLore [Zhao et al., 2024] 16.89 25.52 27.85 14.74 26.03 36.00
+Online Subspace Descent Liang et al. [2024] 16.61 25.86 28.76 18.45 32.57 33.10
+LDAdam [Robert et al., 2025] 17.18 25.94 28.03 18.45 32.57 OOM
+Fira [Chen et al., 2025] 16.39 24.99 27.33 14.68 25.62 47.84
+SubTrack++ (Ours) 16.40 25.06 27.42 15.22 25.54 49.82
+Additionally, Table 9 presents the wall-time consumed by each baseline across all model architectures during pre-training.
+Each run is configured to include exactly 10 subspace updates for SubTrack++ and other baselines employing periodic subspace updates.
+Specifically, models ranging from 60M to 3B use a subspace update interval of 200, resulting in 2,000 total iterations, while the 7B models use an interval of 500, yielding 5,000 iterations.
+Experiments for the 60M to 3B models are conducted on an NVIDIA A100 GPU, while the 7B model experiments are run on an NVIDIA RTX A6000.
+Table 9: Wall-time comparison for pre-training LLaMA-based architectures of varying sizes on the C4 dataset.
+The number of iterations is set to ensure 10 subspace updates for methods using periodic subspace updates: 2k for models ranging from 60M to 3B, and 5k for the 7B model.
+SubTrack++ achieves the lowest wall-time among all baselines that support full-parameter training on large models.
+The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with .
+Since this can impact wall-time comparisons, it is more appropriate to compare runs within the same cluster.
+60M 130M 350M 1B 3B 7B
+r=128 r=256 r=256 r=512 r=512 r=1024
+Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 12726.2
+BAdam [Luo et al., 2024] 511.3 779.2 961.6 798.6 1004.1 7283.7
+GaLore [Zhao et al., 2024] 547.8 1094.2 1589.0 1729.5 2715.5 21590.4
+Online Subspace Descent [Liang et al., 2024] 662.8 1228.2 1818.6 1438.7 1676.9 21590.4
+LDAdam [Robert et al., 2025] 639.9 1342.2 2083.4 2780.9 4625.4 OOM
+Fira [Chen et al., 2025] 635.3 1180.8 1729.7 1938.5 2898.4 22554.3
+SubTrack++ (Ours) 627.6 1140.6 1593.2 1304.3 1517.6 16491.7
+Appendix G Pre-Training LLama-Based Architectures
+We pre-trained all six Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 10.
+Hyperparameters are selected based on reported values in GaLore paper [Zhao et al., 2024].
 Table 10: Hyperparameters of pre-training Llama-based architectures.
-60M 130M 350M 1B 3B
-Architectural Hidden 512 768 1024 2048 4096
-Parameters Intermediate 1376 2048 2736 5461 11008
-Heads 8 12 16 24 32
-Layers 8 12 24 32 32
-Shared Parameters Learning Rate 1e-3 1e-3 1e-3 1e-5 1e-5
-Batch Size 128 128 64 8 8
+60M 130M 350M 1B 3B 7B
+Architectural Hidden 512 768 1024 2048 2560 4096
+Parameters Intermediate 1376 2048 2736 5461 6848 11008
+Heads 8 12 16 24 32 32
+Layers 8 12 24 32 32 32
+Shared Parameters Learning Rate 1e-3 1e-3 1e-3 1e-4 1e-4 1e-4
+Batch Size 128 128 64 8 8 4
+Gradient Accumulation 2 2 2 2 2 4
 Iterations 10K
-Gradient Accumulation 2
 Gradient Clipping 1.0
@@
 dtype bfloat16
-SubTrack & Rank 128 256 256 512 512
-GaLore Parameters Subspace Update Interval 200
-SubTrack-Grad Step-Size 10
+Low-Rank Optimizer Rank 128 256 256 512 512 1024
+Methods Parameters Subspace Update Interval 200 200 200 200 200 500
+SubTrack++ Step-Size 10
 BAdam Parameters Block Swithch Interval 100
 Switch Mode Random
-Appendix G Time and Performance Consistency
-Table 15 demonstrates that increasing the update frequency significantly increases GaLore’s runtime, while the performance of both methods remains comparable.
-This underscores SubTrack-Grad’s efficiency in reducing runtime without sacrificing performance.
-Notably, this advantage becomes even more pronounced for tasks that demand more frequent updates.
-Table 15: Comparing changes in performance while decreasing the subspace update interval (increasing the frequency of subspace updates) on fine-tuning RoBERTa-Base on COLA task.
-500 450 400 350 300 250 200 150 100 50
-GaLore’s Wall-Time 458.79 460.19 471.36 483.04 495.83 520.28 556.14 605.1 715.87 1038.79
-GaLore’s Performance 0.5358 0.5285 0.5385 0.5356 0.5309 0.54821 0.5332 0.5387 0.5463 0.5463
-SubTrack-Grad’s Wall-Time 408.22 408.44 408.47 409.25 409.17 409.76 410.29 411.77 413.35 417.07
-SubTrack-Grad’s Performance 0.5332 0.5382 0.5364 0.5385 0.5354 0.5364 0.5416 0.5385 0.5364 0.5673
-Appendix H Convergence vs. Wall-Time
-Figure 5 depicts the changes in the training loss function relative to wall-time.
-The results demonstrate that SubTrack-Grad’s wall-time reduction has minimal impact on the learning process, showcasing a convergence pattern comparable to other methods without introducing significant computational overhead.
-Figure 5: The figures present changes in training loss relative to wall-time for pre-training the Llama-3B architecture on the C4 dataset and fine-tuning RoBERTa-Large on the BoolQ dataset.
-They demonstrate that SubTrack maintains the overall learning process for both pre-training and fine-tuning while avoiding significant computational overhead.
+Appendix H Subspace Tracking vs. SVD
+Replacing SVD with Grassmannian-based subspace tracking, as opposed to the approach used in GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025], offers two key advantages: (1) Robustness: As shown in Figure 5, Grassmannian tracking better handles noise and abrupt shifts, preserving subspace quality under dynamic conditions.
+(2) Efficiency: It significantly reduces computational overhead (see Table 2 and Appendix D).
+Figure 6 illustrates the wall-time of the experiments in the ablation in Figure 3, highlighting substantial speedups.
+As a result, SubTrack++ achieves SOTA-level performance with GaLore-like memory use, while being one of the fastest methods; clearly demonstrated in Figure LABEL:fig:modular and 6.
+Figure 6: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ .
+Grassmannian tracking exceeds GaLore’s convergence, and significantly reduces wall-time, as also shown in Figure LABEL:fig:modular.

### v3

@@
 Yet, true democratization of LLMs requires simultaneous progress across all three dimensions.
-To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to changes in the optimization subspace.
+To this end, we propose SubTrack++ that leverages Grassmannian gradient subspace tracking combined with projection-aware optimizers, enabling Adam’s internal statistics to adapt to subspace changes.
 Additionally, employing recovery scaling, a technique that restores information lost through low-rank projections, further enhances model performance.
-Our method demonstrates SOTA convergence by exploiting Grassmannian geometry and achieves lowest evaluation loss, outperforming the current SOTA while reducing pre-training wall time by 43% and maintaining the memory footprint on a 1B-parameter Llama model.
+Our method demonstrates SOTA convergence by exploiting Grassmannian geometry, reducing pre-training wall-time by up to 65% and fine-tuning time by 36% compared to existing SOTA methods, while maintaining the same memory footprint.
+Code is at https://github.com/criticalml-uw/SubTrack.
 1 Introduction
-Large Language Models (LLMs) have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity.
+LLMs have demonstrated state-of-the-art performance across a wide range of tasks and are rapidly growing in popularity.
 However, training and fine-tuning these models require significant resources, including extensive hardware and time, which limits their practicality for many applications and increases their environmental impact and carbon footprint.
 (Zhao et al., 2024; Jaiswal et al., 2024; Muhamed et al., 2024; Miles et al., 2024; Modoranu et al., 2024; Hao et al., 2024; Li et al., 2024).
-Several techniques have been proposed to mitigate memory bottlenecks, such as gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020).
+Several techniques have been proposed to mitigate memory bottlenecks(Chen et al., 2016; Rajbhandari et al., 2020).
 LoRA (Hu et al., 2021) and other low-rank adaptation methods (Dettmers et al., 2024; Hu et al., 2021; Yaras et al., 2024; Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Miles et al., 2024) have gained popularity by optimizing a reduced set of parameters.
-Such approaches often assume that the full parameter space is inherently low-rank—an assumption that does not always hold and can lead to suboptimal performance.
-In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of partial parameter tuning and reduced accuracy.
+Such approaches often assume a low-rank parameter space, which can lead to suboptimal performance.
+In addition, methods like BAdam (Luo et al., 2024) and Block-LLM (Ramesh et al., 2024), utilize block coordinate descent to optimize parameter subsets, achieving memory savings at the cost of reduced accuracy.
 Figure 1: We compare baselines on pre-training a 1B-parameter model.
@@
 Overall, SubTrack++ outperforms all baselines in evaluation loss while matching or exceeding them in memory and runtime efficiency.
-However, memory requirements extend beyond trainable parameters, with a significant portion consumed by optimizers for storing element-wise states (Zhao et al., 2024).
-To address this, recent efforts have focused on reducing the optimizer parameters while still targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024; Rajabi and Rambhatla, 2024a).
-Leveraging the fact that gradients evolve in a low-dimensional space during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating the rank-$r$ approximation via singular value decomposition (SVD)—an approach also used in more recent method, Fira (Chen et al., 2025).
-While SVD offers optimal low-rank approximation (Robert et al., 2025), performing it repeatedly can pose several challenges.
-First, it is computationally intensive, and alternatives like random projections (as in APOLLO (Zhu et al., 2025) and GoLore (He et al., 2025)) or approximation methods for estimating dominant singular values (as in LDAdam (Robert et al., 2025), Online Subspace Descent (Liang et al., 2024) and Rajabi and Rambhatla (2024b)) match or outperform SVD in practice.
+However, memory requirements extend beyond trainable parameters, with a significant portion consumed by the optimizer’s states (Zhao et al., 2024).
+Recent efforts have focused on reducing this space while targeting full parameter training (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022; Zhang et al., 2024; Modoranu et al., 2024; Zhao et al., 2024; Muhamed et al., 2024).
+Leveraging the low-dimensional nature of gradients during gradient descent (Gur-Ari et al., 2018; Schneider et al., 2024; Yaras et al., 2023) , GaLore (Zhao et al., 2024) reduces memory usage by projecting gradients into a low-rank subspace and periodically updating this approximation via singular value decomposition (SVD).
+While SVD offers optimal low-rank approximation (Robert et al., 2025), it can pose several challenges.
+First, it is computationally intensive, and alternatives which use random projections (Zhu et al., 2025) or approximation methods to estimate dominant singular values (Robert et al., 2025; Liang et al., 2024), match or outperform SVD in practice.
 Moreover, SVD is sensitive to noise (Vaswani et al., 2018; He et al., 2025) and tends to degrade in late training stages when gradients are small, often hindering convergence (He et al., 2025).
-Geometry-based methods have demonstrated strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Rajabi and Rambhatla, 2024c; Blocker et al., 2023).
-In this context, using Grassmannian manifolds for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016).
-Grassmannian is the set of $r$-dimensional subspaces in a d-dimensional space, and the ultimate goal of low-rank optimization is to find the proper subspace in this space.
-This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017), outperforming online or robust Principle Component Analysis (PCA) techniques (Li, 2004; Balzano et al., 2018; Chakraborty et al., 2017).
+Geometry-based methods have shown strong performance in various machine learning applications (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023); Grassmannian is the manifold of all subspaces of dimensions $r$ in a space of dimensions $d$, and using Grassmannian for subspace tracking has led to structurally embedded information, lower computational complexity, and improved performance (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Blocker et al., 2023; Chakraborty et al., 2017; Zhang and Balzano, 2016).
+This line of work has demonstrated robustness and efficiency in high-dimensional, noisy environments (Zhang et al., 2018; Balzano et al., 2011; He et al., 2011; Balzano et al., 2018; Chakraborty et al., 2017).
 Their natural robustness against perturbations and strong theoretical guarantees make them particularly well-suited to tracking the evolving gradient subspaces encountered in LLM training.
-(Balzano, 2012; Balzano et al., 2011).
-Figure 2: Visualization of Grassmannian-based subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace.
-The tangent vector $\nabla F$ is computed via the derivative of a loss function that measures the subspace estimation error.
-The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$, on the Grassmannian manifold to minimize estimation error.
-To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient method for tracking the gradient subspace.
-This way, instead of reconstructing low-rank approximations via expensive SVD or random projections, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection.
+Figure 2: Visualization of Grassmannian subspace tracking: Between subspace updates, gradients are projected onto a fixed subspace.
+The tangent vector $\nabla F$ is computed via the derivative of a loss function, measuring the subspace estimation error.
+The subspace is then updated by moving along the corresponding geodesic, determined by $\nabla F$ to minimize estimation error.
+To this end, we employ subspace tracking on Grassmannian geodesics (Figure 2), to develop a geometry-based, time- and memory-efficient training method.
+This way, instead of reconstructing low-rank approximations via expensive SVD, we can efficiently leverage previously computed subspaces and the estimation error to better adjust the projection.
 We also incorporate subspace shifts into Adam’s first and second momentum update rules, ensuring proper alignment with coordinate changes using a Projection-Aware Optimizer (Robert et al., 2025).
-Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection to boost performance by utilizing the scaling information of the low-rank optimizer Chen et al.
+Additionally, we recover and scale (Recovery Scaling) the gradient information lost during low-rank projection by utilizing the scaling information of the low-rank, state-full optimizer Chen et al.
 (2025); Zhu et al.
 (2025).
-To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024), FiRA (Chen et al., 2025), LDAdam (Robert et al., 2025)), while maintaining GaLore’s memory footprint.
-It outperforms online PCA methods like Online Subspace Descent (Liang et al., 2024) in subspace tracking and achieves SOTA convergence and evaluation loss across all strong baselines; Comparison provided in Figure 1.
+To summarize, SubTrack++ is a projection-aware geometry-based approach that supports full-parameter training and incorporates recovery scaling, offering superior time efficiency compared to SVD or PowerSGD methods (e.g., GaLore (Zhao et al., 2024; Chen et al., 2025; Robert et al., 2025), while maintaining GaLore’s memory footprint.
+It also outperforms online PCA subspace tracking methods (Liang et al., 2024) and achieves SOTA convergence and evaluation loss across all strong baselines; with comparison provided in Figure 1 and Figure 7.
 2 SubTrack++
 Challenges of Low-Rank Optimization.
-Projecting gradients into a low-rank subspace reduces memory overhead and enables scalable LLM training, but it introduces important trade-offs.
-First, gradient subspaces can be unstable and require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive.
+Projecting gradients into a low-rank subspace reduces memory footprint and enables scalable LLM training, but it introduces important trade-offs.
+First, gradient subspaces require adaptive tracking; while SVD-based methods can capture these shifts (Zhao et al., 2024; Chen et al., 2025), they are computationally expensive.
 To address this, recent work has explored cheaper approximations and random projections (Robert et al., 2025; Liang et al., 2024; Zhu et al., 2025; He et al., 2025).
-Furthermore, optimizers like Adam assume a fixed coordinate system, so any changes in the subspace must be reflected in their internal states to preserve consistent momentum updates.
-Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can lead to improved performance (Chen et al., 2025; Zhu et al., 2025).
+Furthermore, optimizers like Adam assume a fixed coordinate system, and subspace changes must be reflected in their internal states for a consistent momentum updates.
+Finally, low-rank projections inherently discard some gradient components, that recovering and utilizing these discarded signals can boost performance (Chen et al., 2025; Zhu et al., 2025).
 Overview.
-We propose SubTrack++ , a memory- and time-efficient method for full, simultaneous parameter tuning that embeds geometric insights into low-rank optimization.
-It improves efficiency and performance via three core components: 1) Subspace Tracking—Grassmannian-based updates that refine projections using estimation error and subspace history; 2) Projection-Aware Optimizer—adjusting Adam’s internal states to account for shifting subspaces; and 3) Recovery Scaling—rescaling discarded gradient components to reclaim lost information.
+SubTrack++ , a memory- and time-efficient method, embeds geometric insights into low-rank optimization, and improves efficiency and performance via three core components: 1) Grassmannian subspace tracking that refines projections via estimation error and subspace history; 2) projection-aware optimizer which adapts Adam’s state to account for evolving subspaces; and 3) recovery scaling that restores lost information by scaling discarded gradient components.
 Subspace Tracking.
-In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian—the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024)—which is the formal definition and goal of low-rank projections.
-This geometric perspective offers three key benefits: 1) By moving along Grassmann geodesics, it refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in (Zhao et al., 2024; Chen et al., 2025; He et al., 2025).
+In SubTrack++ , subspace estimation is framed as selecting a point on the Grassmannian, the manifold of all $d$-dimensional subspaces in an $n$-dimensional space (Bendokat et al., 2024).
+This geometric perspective offers three key benefits: 1) It refines the subspace using prior subspace and estimation error, avoiding full reinitialization as done in Zhao et al.
+(2024); Chen et al.
+(2025); He et al.
+(2025).
 2) Updates rely on lightweight algebraic operations, minimizing time and memory costs.
 3) Controlled subspace shifts improve robustness against noise and abrupt changes.
-The initial subspace is computed using SVD as shown in (1). $G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 00; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
+The initial subspace is computed using SVD as shown in (1).
+$G_{0}\in\mathbb{R}^{m\times n}$ is the gradient matrix at step 0; $U$, $S$, and $V$ are its SVD components, and $r$ is the specified rank.
 $\small G_{0}=USV^{\top}\approx\sum_{i=1}^{r}s_{i}u_{i}v_{i}^{\top}$ (1)
@@
 The gradient is projected as $\widetilde{G}_{t}=S_{t}^{\top}G_{t}\in\mathbb{R}^{r\times n}$, and the optimizer operates in this reduced space, significantly lowering memory and state overhead.
-To track subspace drift, we adjust the core subspace every $k$ steps, defined as the subspace update interval, by minimizing its Euclidean distance to the current gradient.
-The objective is formalized in the cost function (2).
+To account for subspace drift, the core subspace is updated every $k$ steps (the subspace update interval) by minimizing the cost function (2), which measures its Euclidean distance to the current gradient.
 $\small F(S_{t})=\min_{A}\|S_{t}A-G_{t}\|^{2}_{F},$ (2)
@@
 $\small\nabla F=(I-S_{t}S_{t}^{\top})\frac{\partial F}{\partial S_{t}}=\frac{\partial F}{\partial S_{t}}=-2RA^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ (4)
-The tangent vector $\nabla F$ provides the direction for adjusting the subspace by accounting for the error lying in the orthogonal complement.
-However, to minimize changes to the subspace, SubTrack++ first computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, and represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ in the final equality of (4).
-This approximation is then used to update the subspace.
+For optimizing the loss function (2), the subspace should be moved in the direction of $\nabla F$ to reduce the estimation error.
+However, to control subspace changes, SubTrack++ computes a rank-$1$ approximation of $\nabla F$, determined by its largest singular value and the corresponding singular vector obtained from its SVD, represented as $\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$.
+This approximation is then used for subspace update.
 As shown by Edelman et al.
 (1998); Bendokat et al.
-(2024), from which we have included the necessary definitions and theorem in section 3, we can move along a Grassmannian geodesic guided by the computed tangent vector, taking a step of size $\eta$, as presented in (5).
-$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$ (5)
+(2024), we can move along a Grassmannian geodesic guided by rank-1 estimation of $\nabla F$, with a step-size $\eta$, as presented in (5).
+$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ -\sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$ (5)
 This update rule preserves the orthonormality of $S_{t+1}$, ensuring it remains on the Grassmannian.
@@
   if $t$ mod $k==0$ then
-        $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$ $S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
+   $G_{lr}=\arg\min_{A}\|(S_{t-1}A-G_{t})\|^{2}$, and $R=G_{t}-S_{t-1}G_{lr}$ $\nabla F=-2RG_{lr}^{\top}\approx\widehat{U}_{F}\widehat{\Sigma}_{F}\widehat{V}^{\top}_{F}$$S_{t}=(S_{t-1}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ -\sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t-1}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
    $M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$ $\widetilde{G}_{t}=S_{t}^{\top}G_{t}$: low-rank projection of $G_{t}$} $\mathcal{V}_{t}\leftarrow\beta_{2}\cdot[(1-\beta_{2}^{t-1})|(S_{t}^{\top}S_{t-1})^{2}\cdot(\mathcal{V}_{t-1}-M_{t-1}^{2})+(S_{t}^{\top}S_{t-1}\cdot M_{t-1})^{2}|]+(1-\beta_{2})\cdot\widetilde{G}_{t}^{2}$
@@
   if $\frac{\Lambda_{t}}{\Lambda_{t-1}}>\zeta$ then $\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$
-     $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $-\alpha\cdot\Lambda_{t}$
+  $W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}$ $\alpha\cdot\Lambda_{t}$
 end for
 Projection-Aware Optimizer.
-In Adam, the first and second momentum update rule is shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
+In Adam, the first and second momentum update rules are as shown in (6) and (7), respectively (reminder: $\widetilde{G}_{t}$ is the projection of gradient into low-rank subspace).
 $\small M_{t}\leftarrow\beta_{1}\cdot M_{t-1}+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (6)
@@
 To address this, at each subspace update step, we modify Adam’s original update rules in (6) and (7), replacing them with projection-aware counterparts represented in (8) and (9), which reflect subspace changes into optimizer statistics (Robert et al., 2025).
-Further details regarding this projections can be found in Appendix C.
+Further details regarding these projections can be found in Appendix C.
 $\small M_{t}\leftarrow\beta_{1}\cdot(S_{t}^{\top}S_{t-1}M_{t-1})+(1-\beta_{1})\cdot\widetilde{G}_{t}$ (8)
@@
 $\small W_{t}\leftarrow W_{t-1}-\alpha\cdot\widehat{G}_{t}-\alpha\cdot\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ (10)
-Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
+Here $\alpha$ is the learning rate and $\phi_{t}(G_{t})$ is the column-wise scaling factor computed based on the low-rank gradient representation, $\widetilde{G}_{t}$, and the optimizer’s processed output, $\widetilde{G}_{t}^{O}$ as:
 $\small\phi_{t}(G_{t})_{i}=\frac{\|\widetilde{G}_{t,:,i}^{O}\|}{\|\widetilde{G}_{t,:,i}\|}$ (11)
-As suggested by Fira (Chen et al., 2025), this additional term must be controlled using a gradient-clipping-like method to ensure training stability.
-Building on their observation, we also limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
+Following Fira’s observation (Chen et al., 2025), we employ a gradient-clipping-inspired mechanism to stabilize training.
+Specifically, we limit the growth rate of $\Lambda_{t}=\phi_{t}(G_{t})(G_{t}-S_{t}\widetilde{G}_{t})$ by a factor $\zeta$, and apply a correction whenever it exceeds this threshold as:
 $\small\Lambda_{t}\leftarrow\frac{\Lambda_{t}}{\|\Lambda_{t}\|}\cdot\zeta\|\Lambda_{t-1}\|$ (12)
@@
 The proof of Theorem 3 can be found in Appendix B. Leveraging this theorem and our notation in section 2, one can easily verify that the subspace update rule is as follows:
-$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ \sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
+$\small S_{t+1}(\eta)=(S_{t}\widehat{V}_{F}\quad\widehat{U}_{F})\begin{pmatrix}\cos{\widehat{\Sigma}_{F}\eta}\\ -\sin{\widehat{\Sigma}_{F}\eta}\end{pmatrix}\widehat{V}^{\top}_{F}+S_{t}(I-\widehat{V}_{F}\widehat{V}^{\top}_{F})$
 This update rule generally converges to a stable subspace if the step size $\eta$ decreases over time (Balzano et al., 2011).
@@
 4 Experiments and Results
-To ensure fair comparison, we evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
-Pre-Training Experiments.
-We pre-trained several Llama-based models on the C4 dataset, with results in Table 1.
-To ensure a fair comparison, we benchmarked against a diverse set of baselines.
-BAdam (Luo et al., 2024) is notably efficient in both memory and time, but its reliance on partial parameter updates leads to reduced performance.
-GaLore (Zhao et al., 2024) supports full-parameter training with low memory usage, though this comes at the cost of more than 150% increase in wall-time relative to AdamW in large models.
-Table 1: We compare evaluation loss for pre-training Llama-based architectures on the C4 dataset over 10K iterations.
+We evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
+Table 1: We compare evaluation loss ($\downarrow$) for pre-training Llama-based architectures on the C4 dataset over 10k iterations.
 SubTrack++ outperforms all other baselines in nearly every configuration.
@@
 Fira (Chen et al., 2025) 3.80 3.55 3.56 6.31 6.50 6.83
-SubTrack++ (Ours) 3.43 3.24 3.29 4.51 4.49 4.62
-As another relative baseline, Online Subspace Descent (Liang et al., 2024) uses Online PCA to track gradient subspace.
-LDAdam (Robert et al., 2025) and Fira (Chen et al., 2025) deliver strong performance at the cost of increasing wall-time even more than GaLore.
-Across these methods, SubTrack++ achieves SOTA results while using 14% less memory and running 43% faster than LDAdam, the most accurate baseline on 1B model.
-Hyperparameter details are provided in Appendix G, with a detailed wall-time and memory analysis in Appendix F.
+SubTrack++ (Ours) 3.43 3.24 3.29 4.52 4.50 4.63
+Pre-Training Experiments.
+We pre-trained several Llama-based models on the C4 dataset, with results in Table 1.
+To ensure a fair comparison, we benchmarked against a diverse set of baselines.
+While all compared methods aim for memory-efficient training, their architectural principles yield distinct computational and convergence trade-offs.
+BAdam (Luo et al., 2024) achieves strong memory and time efficiency, yet its partial parameter update strategy compromises final performance.
+GaLore (Zhao et al., 2024) constrains gradients to a low-rank subspace estimated via periodic SVD, an approach sensitive to noise and costly to compute, while discarding information residing in the orthogonal complement.
+Fira (Chen et al., 2025) introduces norm-based recovery scaling to mitigate this information loss, but its reliance on frequent SVD still leads to substantial wall-time overhead.
+LDAdam (Robert et al., 2025) replaces SVD with PowerSGD-based iterative updates and a projection-aware optimizer that synchronizes internal states with evolving subspaces, improving convergence and stability but incurring high per-step costs due to continual subspace updates.
+It also adds an extended error-feedback mechanism to compensate for both gradient and optimizer compression to target GaLore’s shortcomings.
+Online Subspace Descent (OSD) (Liang et al., 2024), our tracking-based baseline, further reduces complexity by employing Online PCA for subspace tracking.
+SubTrack++ departs from these formulations by treating subspace evolution as a geometric tracking problem on the Grassmannian.
+It performs efficient rank-1 geodesic updates that reuse historical subspace information, inherently preserving consistency and avoiding the instability of discrete subspace resets.
+Simultaneously, it aligns Adam’s internal states through projection-aware optimizer and leverages recovery scaling to reintegrate the lost information.
+This unified approach yields a new training paradigm that retains the memory efficiency of low-rank methods, while achieving on-par runtime efficiency compared to fastest baselines like APOLLO (Zhu et al., 2025).
+SubTrack++ sets new state-of-the-art results across model scales, running 43% faster than LDAdam, the strongest prior baseline, on the 1B model, and 67% faster on the 3B model (see Table 6).
+Figure LABEL:fig:modular demonstrates the pre-training of a 1B-parameter Llama model across several baselines.
+SubTrack++ has the fastest convergence in both training steps and wall-time, highlighting the effectiveness of geometry-aware optimization in improving performance and reducing resource consumption.
+To further assess its generalization in longer training regimes and larger models, we extend training to 100k steps and compare SubTrack++ with GaLore.
+As shown in Figure LABEL:fig:7b-long, SubTrack++ converges substantially faster, achieving an evaluation loss of 3.37 (a significant improvement over the 10k-step setting), while GaLore reaches 4.64 under the same conditions.
+As shown in Table 1, SubTrack++ occasionally outperforms full-rank training.
+This effect may be attributed to the implicit regularization introduced by low-rank projections, which can enhance generalization in overparameterized models.
+Similar trends have been reported in other studies (Robert et al., 2025; Zhu et al., 2025; Chen et al., 2025).
+In addition, to further validate that convergence on the projected gradient reflects convergence of the original full-gradient, we measured both norms during Llama-1B pre-training.
+The full-gradient norm drops from 0.46 → 0.08, while the projected-gradient norm drops from 0.45 → 0.05, following nearly identical trajectories.
+This confirms that the optimization progress observed on the projected gradient accurately mirrors convergence in the original gradient space, consistent with prior low-rank gradient findings.
+Hyperparameters of pre-training experiments are provided in Appendix E, with detailed runtime and memory reports in Appendix F.
+(a) Loss vs. Training Steps.
+(b) Loss vs. Relative Wall-Time.
 Figure 3: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ .
-While Grassmannian tracking alone matches GaLore’s step-wise convergence, it significantly reduces wall-time, as shown in Figure LABEL:fig:modular.
-Figure LABEL:fig:modular compares pre-training performance of a 1B-parameter Llama model across several baselines.
-SubTrack++ converges fastest in both steps and wall-time, emphasizing the power of geometry-aware optimization in boosting performance, reducing resource demands, and democratizing access to SOTA training methods.
-Ablation Study.
-We conducted an ablation study to evaluate the individual and combined effects of projection-aware optimizers and recovery scaling, integrated with Grassmannian subspace tracking on the 1B Llama model.
-We use the experimental settings outlined in Table 10.
-As shown in Figure 3, projection-aware optimizer and recovery scaling independently yield substantial improvements over pure Grassmannian subspace tracking, reducing loss from 6.53 to 5.43 and 5.28, respectively; while their combination, SubTrack++ , achieves 4.51 surpassing all baselines.
-Notably, these gains are achieved with minimal increases in runtime and memory.
+While Grassmannian tracking alone almost matches GaLore’s step-wise convergence (a), it significantly reduces wall-time (b).
+Ablation Studies.
+We conducted an ablation study to assess the individual and combined contributions of the projection-aware optimizer and recovery scaling, integrated with Grassmannian subspace tracking, on the 1B Llama model.
+Experimental settings are summarized in Table 4.
+As illustrated in Figure 3-b, Grassmannian subspace tracking alone substantially reduces wall-time compared to frequent SVD updates.
+Both the projection-aware optimizer and recovery scaling independently provide notable performance gains over baseline subspace tracking, lowering the loss from 6.53 to 5.43 and 5.28, respectively.
+Their combination, SubTrack++ , further improves the loss to 4.51, surpassing all baselines.
+Importantly, these improvements are achieved with only minimal increases in runtime and memory overhead, thanks to the efficiency of Grassmannian subspace tracking.
+In addition, we conducted ablations on the subspace update rank and update frequency.
+As shown in Figure 4-a, more frequent updates can further improve performance, and the computational efficiency of SubTrack++ enables higher update frequencies with minimal overhead.
+However, excessively frequent updates may impede convergence.
+Figure 4-b shows that rank-1 updates achieve the best performance among all tested values.
+This suggests that making controlled, small adjustments to the underlying subspace helps maintain stability, while updating the subspace along the most informative direction prevents stagnation in a low-rank region and enhances generalization.
 Fine-Tuning Experiments.
-RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE and SuperGLUE tasks; with the results presented in Table 4 and 5, respectively.
-More details and hyperparameters are provided in Appendix E.
-Time and Space Complexity.
-Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$.
-GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration.
-In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira.
-Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods.
-A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D.
-Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.
+RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE (Wang et al., 2019) and SuperGLUE (Sarlin et al., 2020) tasks; with the results presented in Table 7 and 8, respectively.
+We also conducted supervised fine-tuning of the Llama-2-7B-chat-hf model for one epoch on the Alpaca (Taori et al., 2023) dataset.
+In this experiment, SubTrack++ achieved 36% lower wall-time and better performance compared to GaLore.
+The results are presented in Table 9.
+More details and hyperparameters are provided in Appendix G.
+Time and Space Complexity. Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$.
 Table 2: The optimizer’s state parameter count and subspace update time complexity across baselines, given a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$.
@@
 SubTrack++ $mr+2nr$ $O(mnr)$
-(a) Ours, SF=1
-(b) GaLore, SF=1
-(c) Ours, SF=3
-(d) GaLore, SF=3
-Figure 5: Comparison of Grassmannian subspace tracking (Ours) (a, c) and GaLore’s SVD (b, d) on the Ackley Function over 100 optimization steps, with a subspace update interval of 10.
-SF stands for scale factor; with a scale factor of 1, GaLore fails to reach the global minimum due to abrupt jumps.
-At a scale factor of 3, while the minimum is reached, the jump length increases.
-This demonstrates SVD’s sensitivity to noise and abrupt changes, highlighting the robustness of our Subspace Tracking method with its controlled subspace updates.
+GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration.
+In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira.
+Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods.
+A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D.
+Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.
+(a) Loss vs. Subspace Update Frequency.
+(b) Loss vs. Subspace Update Rank.
+Figure 4: Ablation results on (a) update frequency: decreasing the update interval (i.e., increasing the frequency) improves evaluation performance up to a point, but overly frequent updates hinder training convergence.
+(b) update rank: increasing the rank of updates degrades model performance, and beyond a certain threshold, can prevent convergence.
+These results emphasize the importance of controlled subspace adjustments.
 Robust Subspace Tracking.
@@
 GaLore struggles to reach the global minimum within 100 steps at scale factor 1, and although increasing the scale factor to 3 improves performance, it amplifies jumps that hinder convergence in non-convex settings, revealing sensitivity to hyperparameters, noise, and abrupt changes.
+To empirically measure the robustness of SubTrack++ in tracking evolving subspaces, we quantified subspace drift using the norm of the tangent vector $\nabla F$, which reflects the deviation between the projected gradient and the original gradient matrix.
+On Llama-350M pre-training, this norm rapidly decays from 0.06 to below 0.0002 within 5-7 subspace updates and remains near zero thereafter, indicating highly stable and well-aligned subspace tracking.
+This quantitative stability supports our qualitative observations in Figure 6 and aligns with prior findings in Grassmannian optimization.
 5 Related Works
@@
 (2024) propose compressing intermediate activations and reconstructing them during backpropagation to enhance memory efficiency.
+Yen et al.
+(2025) propose adjustments to the LoRA factorization that promotes balanced training and correspondingly update the optimizer’s internal states (i.e., first and second moments) to remain consistent under the change of basis.
 Additionally, Hao et al.
@@
 BAdam (Luo et al., 2024) leverages the block coordinate descent framework to reduce memory consumption while maintaining capabilities comparable to Adam.
+(a) Ours, SF=1
+(b) GaLore, SF=1
+(c) Ours, SF=3
+(d) GaLore, SF=3
+Figure 6: Comparison of Grassmannian subspace tracking (Ours) (a, c) and GaLore’s SVD (b, d) on the Ackley Function over 100 optimization steps, with a subspace update interval of 10.
+SF stands for scale factor; with a scale factor of 1, GaLore fails to reach the global minimum due to abrupt jumps.
+At a scale factor of 3, while the minimum is reached, the jump length increases.
+This demonstrates SVD’s sensitivity to noise and abrupt changes, highlighting the robustness of our subspace tracking method with its controlled subspace updates.
 Gradient Low-Rank Projection.
@@
 GaLore (Zhao et al., 2024) leverages this fact to reduce the optimizer’s memory by projecting gradients into a low-rank subspace and then projecting them back for full parameter tuning.
-This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024; Rajabi and Rambhatla, 2024b).
+This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024).
 However, not all layers’ gradients evolve within a stable low-rank subspace.
 Jaiswal et al.
-(2024); Rajabi and Rambhatla (2024d) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates.
+(2024) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates.
 Grass (Muhamed et al., 2024) reduces memory usage by applying sparse projection matrices to the gradient.
@@
 LDAdam (Robert et al., 2025) performs optimization within lower-dimensional subspaces, incorporating a projection-aware optimization update rule and a generalized error feedback mechanism.
-APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection.
+Projection-Aware APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection.
 Also, Liang et al.
-(2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD, a similar approach is also proposed in Rajabi and Rambhatla (2024a).
-Grassmannian and Subspace Tracking.
+(2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD.
+Geometric Subspace Updates.
 A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve.
 Balzano et al.
-(2011); Rajabi and Rambhatla (2024c) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed.
+(2011) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed.
 Zhang and Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces.
@@
 (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively.
+Mo et al.
+(2025) propose LORO, a low-rank pretraining method that performs Riemannian optimization on the manifold of fixed-rank matrices, enabling parameter- and memory-efficient training by updating low-rank factors via manifold-aware gradients and retractions.
 6 Discussion and Conclusion
@@
 Also benchmarking on models with more than 7B parameters was not feasible regarding our limited time and resources.
+Acknowledgments and Disclosure of Funding
+Sirisha Rambhatla would like to acknowledge support of the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant, RGPIN-2022-03512.
 Appendix A Convergence of SubTrack++
@@
 Let $\otimes$ denote the Kronecker product. Then, $vec(AXB)=(B^{\top}\otimes A)vec(X)$.
-By applying $vec$ to the gradient form given in the theorem, we obtain:
+By applying vecvec to the gradient form given in the theorem, we obtain:
 $g_{t}=vec(G_{t})=vec(\sum_{i}A_{i}+\sum_{i}B_{i}WC_{i})=a_{t}-D_{t}w_{t}$ (16)
@@
 Since both matrices are orthonormal—preserved by the Grassmannian update rule in (5)—the matrix $S_{t}^{\top}S_{t-1}$ represents the change of basis between the two subspaces.
-In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps $t-1$ and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step $t-1$.
+In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps t1t-1 and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step t1t-1.
 $\mathbf{E}_{t,\beta}\left[.\right]$ denotes the exponential time-weighted expectation at time $t$ with decay rate $\beta$.
@@
 This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$.
-The following operation for the partial derivative is $-2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity.
+The following operation for the partial derivative is $2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity.
 The tangent vector computation (4) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$.
@@
 Tangent vector $\Delta F$ $O(m^{2}r)$
-Rank-$1$ approximation of $\Delta F$ $O(mr^{2})$
+Rank-11 approximation of $\Delta F$ $O(mr^{2})$
 Update rule $O(mr^{2})$
@@
 However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack++ is still more efficient with respect to time complexity.
-Appendix E Fine-Tuning Experiments
-As described, we examined SubTrack++ on fine-tuning RoBERTa-base and RoBERTa-large models to evaluate them on GLUE and SuperGLUE benchmarks.
-The results for GLUE task is summarized in Table 4, and SuperGLUE in 5.
-Table 4: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for $r=8$.
-The performance is measured via Accuracy for SST-2 and RTE tasks, F1 for MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA.
-The best results are marked in bold, with the second-best performance underlined.
-COLA STS-B MRPC RTE SST-2
-Full-Rank 62.57 91.03 91.32 77.98 94.27
-BAdam [Luo et al., 2024] 54.44 89.01 91.35 68.59 94.15
-GaLore [Zhao et al., 2024] 58.54 90.61 91.30 74.37 94.50
-LDAdam [Robert et al., 2025] 58.81 90.90 92.22 76.53 94.27
-SubTrack++ (Ours) 58.29 90.90 91.84 76.53 90.83
-Table 5: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with $r=8$.
-The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, and F1 for CB.
-The best results are marked in bold, with the second-best performance underlined.
-BoolQ CB COPA WIC WSC AXg
-Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30
-GaLore [Zhao et al., 2024] 85.44 88.85 80.00 71.47 63.46 100.00
-BAdam [Luo et al., 2024] 82.51 53.28 59.00 70.38 60.58 51.85
-LDAdam Robert et al. [2025] 85.75 56.16 80.00 71.00 64.42 70.37
-SubTrack++ (Ours) 84.68 89.66 88.00 70.06 63.46 100.00
-Table 6: Hyperparameters of fine-tuning RoBERTa-Base on GLUE tasks.
-SST-2 MRPC CoLA RTE STS-B
-Shared Parameters Batch Size 16 16 32 16 16
-# Epochs 30
-Max Seq. Len. 512
-Low-Rank Optimizer Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
-Methods Parameters SubTrack Step-Size 0.01 3.0 5.0 15.0 10.0
-Subspace Update Interval 500
-Rank Config 8
-$\alpha$ 2
-BAdam Parameters Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
-Block Switch Interval 100
-Switch Mode Random
-Table 7: Hyperparameters of fine-tuning RoBERTa-Large on SuperGLUE tasks.
-BoolQ CB COPA WIC WSC AXg
-Shared Parameters Batch Size 16
-# Epochs 30
-Learning Rate 1e-5
-Max Seq. Len. 512
-Low-Rank Optimizer SubTrack++ Step-Size 0.1 1.0 50.0 100.0 1.0 1.0
-Methods Parameters Subspace Update Interval 500 100 100 500 250 100
-Rank Config. 8
-$\alpha$ 4
-BAdam Parameters Block Switch Interval 100 50 50 100 50 50
-Switch Mode Random
-The hyperparameters for fine-tuning RoBERTa-base are detailed in Table 6, matching those reported in the GaLore [Zhao et al., 2024] for rank-$8$ subspaces, with a subspace update interval set at 500 iterations.
-We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al.
-[2024], as detailed in Table 7, with the exception that we fine-tuned each task for 30 epochs.
-Appendix F Memory and Time Comparison
-Table 8 presents the the peak memory consumption measured to compare SubTrack++ and other baselines.
-It shows that SubTrack++ requires on-par or better memory compared to GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025] while getting state-of-the-art results.
-As detailed in Table 2, all baselines except BAdam [Luo et al., 2024] use the same number of optimizer parameters; therefore, any differences in their peak memory consumption stem from variations in their runtime parameters.
-Table 8: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset.
-The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with .
-SubTrack++ demonstrates better or on-par memory compared to other low-rank methods that allows full-parameter training.
-60M 130M 350M 1B 3B 7B
-r=128 r=256 r=256 r=512 r=512 r=1024
-Full-Rank 16.86 25.32 28.67 18.83 34.92 50.50
-BAdam [Luo et al., 2024] 13.34 20.01 16.45 9.18 14.75 22.35
-GaLore [Zhao et al., 2024] 16.89 25.52 27.85 14.74 26.03 36.00
-Online Subspace Descent Liang et al. [2024] 16.61 25.86 28.76 18.45 32.57 33.10
-LDAdam [Robert et al., 2025] 17.18 25.94 28.03 18.45 32.57 OOM
-Fira [Chen et al., 2025] 16.39 24.99 27.33 14.68 25.62 47.84
-SubTrack++ (Ours) 16.40 25.06 27.42 15.22 25.54 49.82
-Additionally, Table 9 presents the wall-time consumed by each baseline across all model architectures during pre-training.
-Each run is configured to include exactly 10 subspace updates for SubTrack++ and other baselines employing periodic subspace updates.
-Specifically, models ranging from 60M to 3B use a subspace update interval of 200, resulting in 2,000 total iterations, while the 7B models use an interval of 500, yielding 5,000 iterations.
-Experiments for the 60M to 3B models are conducted on an NVIDIA A100 GPU, while the 7B model experiments are run on an NVIDIA RTX A6000.
-Table 9: Wall-time comparison for pre-training LLaMA-based architectures of varying sizes on the C4 dataset.
-The number of iterations is set to ensure 10 subspace updates for methods using periodic subspace updates: 2k for models ranging from 60M to 3B, and 5k for the 7B model.
-SubTrack++ achieves the lowest wall-time among all baselines that support full-parameter training on large models.
-The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with .
-Since this can impact wall-time comparisons, it is more appropriate to compare runs within the same cluster.
-60M 130M 350M 1B 3B 7B
-r=128 r=256 r=256 r=512 r=512 r=1024
-Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 12726.2
-BAdam [Luo et al., 2024] 511.3 779.2 961.6 798.6 1004.1 7283.7
-GaLore [Zhao et al., 2024] 547.8 1094.2 1589.0 1729.5 2715.5 21590.4
-Online Subspace Descent [Liang et al., 2024] 662.8 1228.2 1818.6 1438.7 1676.9 21590.4
-LDAdam [Robert et al., 2025] 639.9 1342.2 2083.4 2780.9 4625.4 OOM
-Fira [Chen et al., 2025] 635.3 1180.8 1729.7 1938.5 2898.4 22554.3
-SubTrack++ (Ours) 627.6 1140.6 1593.2 1304.3 1517.6 16491.7
-Appendix G Pre-Training LLama-Based Architectures
-We pre-trained all six Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 10.
-Hyperparameters are selected based on reported values in GaLore paper [Zhao et al., 2024].
+Appendix E Pre-Training Llama-Based Architectures
+We pre-trained all six Llama-based architectures for 10k iterations using hyperparameters reported in Table 4.
+To demonstrate the generalizability of the proposed method, we also present results from pre-training the 7B architecture with both SubTrack++ and GaLore for 100k iterations, as shown in Figure LABEL:fig:7b-long.
+SubTrack++ maintains its advantage in terms of faster convergence and superior performance.
+The hyperparameters of this run are identical to those reported in Table 4, except for the number of iterations which is 100k.
+Also the bar-chart version of Figure 1 is represented in Figure 7.
+Figure 7: We compare baselines on pre-training a 1B-parameter model.
+(a) SubTrack++ achieves the lowest evaluation loss across all methods.
+(b) Its peak memory usage is significantly lower than APOLLO and LDAdam, and on par with GaLore and Fira.
+(c) In terms of wall-time, SubTrack++ incurs minimal overhead relative to APOLLO and is markedly faster than GaLore, Fira, and LDAdam.
+Overall, SubTrack++ outperforms all baselines in evaluation loss while matching or exceeding them in memory and runtime efficiency.
 Table 4: Hyperparameters of pre-training Llama-based architectures.
@@
 Methods Parameters Subspace Update Interval 200 200 200 200 200 500
-SubTrack++ Step-Size 10
+SubTrack++ Step-Size 10000
 BAdam Parameters Block Swithch Interval 100
 Switch Mode Random
-Appendix H Subspace Tracking vs. SVD
-Replacing SVD with Grassmannian-based subspace tracking, as opposed to the approach used in GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025], offers two key advantages: (1) Robustness: As shown in Figure 5, Grassmannian tracking better handles noise and abrupt shifts, preserving subspace quality under dynamic conditions.
-(2) Efficiency: It significantly reduces computational overhead (see Table 2 and Appendix D).
-Figure 6 illustrates the wall-time of the experiments in the ablation in Figure 3, highlighting substantial speedups.
-As a result, SubTrack++ achieves SOTA-level performance with GaLore-like memory use, while being one of the fastest methods; clearly demonstrated in Figure LABEL:fig:modular and 6.
-Figure 6: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ .
-Grassmannian tracking exceeds GaLore’s convergence, and significantly reduces wall-time, as also shown in Figure LABEL:fig:modular.
+Appendix F Memory and Time Comparison
+Table 5 presents the the peak memory consumption measured to compare SubTrack++ and other baselines.
+It shows that SubTrack++ requires on-par or better memory compared to GaLore [Zhao et al., 2024] while getting state-of-the-art results.
+As detailed in Table 2, all baselines except BAdam [Luo et al., 2024] use the same number of optimizer parameters; therefore, any differences in their peak memory consumption stem from variations in their runtime parameters.
+Table 5: Peak memory consumption of pre-training Llama-based architectures on C4 dataset.
+The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with .
+SubTrack++ demonstrates better or on-par memory compared to other low-rank methods that allows full-parameter training.
+60M 130M 350M 1B 3B 7B
+r=128 r=256 r=256 r=512 r=512 r=1024
+Full-Rank 16.86 25.32 28.67 18.83 34.92 50.50
+BAdam [Luo et al., 2024] 13.34 20.01 16.45 9.18 14.75 22.35
+GaLore [Zhao et al., 2024] 16.89 25.52 27.85 14.74 26.03 36.00
+Online Subspace Descent Liang et al. [2024] 16.61 25.86 28.76 18.45 32.57 33.10
+LDAdam [Robert et al., 2025] 17.18 25.94 28.03 18.45 32.57 OOM
+Fira [Chen et al., 2025] 16.39 24.99 27.33 14.68 25.62 47.84
+SubTrack++ (Ours) 16.40 25.06 27.42 15.22 25.54 49.82
+Table 6: Wall-time comparison of pre-training Llama-based architectures on the C4 dataset.
+The number of iterations is set to ensure 10 subspace updates for methods using periodic subspace adjustments.
+SubTrack++ achieves the lowest wall-time among all baselines that support full-parameter training on large models.
+The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with .
+Since this can impact wall-time comparisons, it is more appropriate to compare runs within the same cluster.
+60M 130M 350M 1B 3B 7B
+r=128 r=256 r=256 r=512 r=512 r=1024
+Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 12726.2
+BAdam [Luo et al., 2024] 511.3 779.2 961.6 798.6 1004.1 7283.7
+GaLore [Zhao et al., 2024] 547.8 1094.2 1589.0 1729.5 2715.5 21590.4
+Online Subspace Descent [Liang et al., 2024] 662.8 1228.2 1818.6 1438.7 1676.9 21590.4
+LDAdam [Robert et al., 2025] 639.9 1342.2 2083.4 2780.9 4625.4 OOM
+Fira [Chen et al., 2025] 635.3 1180.8 1729.7 1938.5 2898.4 22554.3
+SubTrack++ (Ours) 627.6 1140.6 1593.2 1304.3 1517.6 16491.7
+Additionally, Table 6 presents the wall-time consumed by each baseline across all model architectures during pre-training.
+Each run is configured to include exactly 10 subspace updates for SubTrack++ and other baselines employing periodic subspace updates.
+Specifically, models ranging from 60M to 3B use a subspace update interval of 200, resulting in 2,000 total iterations, while the 7B models use an interval of 500, yielding 5,000 iterations.
+Experiments for the 60M to 3B models are conducted on an NVIDIA A100 GPU, while the 7B model experiments are run on an NVIDIA RTX A6000.
+Appendix G Fine-Tuning Experiments
+As described, we examined SubTrack++ on fine-tuning RoBERTa-base and RoBERTa-large models to evaluate them on GLUE and SuperGLUE benchmarks.
+The results for GLUE task is summarized in Table 7, and SuperGLUE in 8.
+Table 7: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for $r=8$.
+The performance is measured via Accuracy ($\uparrow$) for SST-2 and RTE tasks, F1 ($\uparrow$) for MRPC, Pearson Correlation ($\uparrow$) for STS-B, and Matthews Correlation ($\uparrow$) for COLA.
+The best results are marked in bold, with the second-best performance underlined.
+COLA STS-B MRPC RTE SST-2
+Full-Rank 62.57 91.03 91.32 77.98 94.27
+BAdam [Luo et al., 2024] 54.44 89.01 91.35 68.59 94.15
+GaLore [Zhao et al., 2024] 58.54 90.61 91.30 74.37 94.50
+LDAdam [Robert et al., 2025] 58.81 90.90 92.22 76.53 94.27
+SubTrack++ (Ours) 58.55 90.95 92.04 78.34 90.02
+Table 8: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with $r=8$.
+The performance is measured via Accuracy ($\uparrow$) for COPA, WIC, WSC, BoolQ, and AXg tasks, and F1 ($\uparrow$) for CB.
+The best results are marked in bold, with the second-best performance underlined.
+BoolQ CB COPA WIC WSC AXg
+Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30
+GaLore [Zhao et al., 2024] 85.44 88.85 80.00 71.47 63.46 100.00
+BAdam [Luo et al., 2024] 82.51 53.28 59.00 70.38 60.58 51.85
+LDAdam Robert et al. [2025] 85.75 56.16 80.00 71.00 64.42 70.37
+SubTrack++ (Ours) 85.38 83.96 82.00 70.70 62.5 100.00
+Table 9: Comparing final evaluation perplexity ($\downarrow$) and wall-time of fine-tuning Llama-2-7B-chat-hf on Alpaca dataset with GaLore and SubTrack++ .
+Method Evaluation Perplexity Wall-Time (min)
+GaLore 2.43 110.8
+SubTrack++ 2.41 73.2
+Table 10: Hyperparameters of fine-tuning RoBERTa-Base on GLUE tasks.
+SST-2 MRPC CoLA RTE STS-B
+Shared Parameters Batch Size 16 16 32 16 16
+# Epochs 30
+Max Seq. Len. 512
+Low-Rank Optimizer Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
+Methods Parameters SubTrack Step-Size 0.1 3.0 5.0 15.0 10.0
+Subspace Update Interval 500
+Rank Config 8
+$\alpha$ 2
+BAdam Parameters Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
+Block Switch Interval 100
+Switch Mode Random
+Table 11: Hyperparameters of fine-tuning RoBERTa-Large on SuperGLUE tasks.
+BoolQ CB COPA WIC WSC AXg
+Shared Parameters Batch Size 16
+# Epochs 30
+Learning Rate 1e-5
+Max Seq. Len. 512
+Low-Rank Optimizer SubTrack++ Step-Size 0.1 10.0 10.0 100.0 1.0 1.0
+Methods Parameters Subspace Update Interval 500 100 100 500 250 100
+Rank Config. 8
+$\alpha$ 4
+BAdam Parameters Block Switch Interval 100 50 50 100 50 50
+Switch Mode Random
+Table 12: Hyperparameters of fine-tuning Llama-2-7B-chat-hf on Alpaca dataset.
+Parameter Value
+Subspace Update Interval 500
+Rank 1024
+$\alpha$ 0.25
+Target Modules att, mlp
+Batch Size 8
+Epoch 1
+The hyperparameters for fine-tuning RoBERTa-base are detailed in Table 10, matching those reported in the GaLore [Zhao et al., 2024] for rank-$8$ subspaces, with a subspace update interval set at 500 iterations.
+We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al.
+[2024], as detailed in Table 11, with the exception that we fine-tuned each task for 30 epochs.
+The results of supervised fine-tuning of the Llama-2-7B-chat-hf model on the Alpaca dataset on an Nvidia-H100 GPU are presented in Table 9.
+The fine-tuning was performed for one epoch, and the corresponding hyperparameters are listed in Table 12.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We included the overview of method, achieved results, and the scope of experiments in abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The limitations and future directions is included in the Discussion section. Also the computational complexity is reported.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: The theoretical proofs are aligned with well-known prior works while showing improvement over them, and backed up by strong sources for Grassmannian manifolds geometry.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: All resources and hyperparameters used for testing the proposed method and baselines are provided in the supplemental material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We will include the code in the main paper after submission, and the anonymized repository in supplemental material of this version.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All the details regarding resources and hyperparameters for the proposed method and baselines are included in supplemental material.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34.
+Justification: As the experiments of these paper are highly resource- and time-consuming, performing statistical test was not possible for this version.
+However, a variety of experiments are included to ensure generalization.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: All the necessary information is concluded in supplemental material.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Yes, it totally conform.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: This is a foundational research, and further investigation for preserving safety is left for future work.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: It is a foundational research and do not poses such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have cited properly.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The code that is included anonymously on supplemental material is documented.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This term does not apply to our research.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This term does not apply to our research.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: It does not apply to our research.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.